### v1

# On the comparison of ultradifferentiable weight function settings

Abstract.
We summarize, revisit and compare different notions of ultradifferentiable settings defined in terms of weight functions. More precisely, we emphasize the difference between the original Beurling-Björck-setting, when measuring the growth of the Fourier-transform, and the more recent approach by Braun, Meise and Taylor, when controlling the growth of the derivatives similarly like in the weight sequence framework. For the equality of both settings the convexity of the defining weight function is required and, somehow, has become a standard assumption. We study the meaning of this convexity condition and its failure in detail. First, we characterize the inclusion relations and hence the equality for Beurling-Björck-type classes in terms of weight functions analogously as it has already been done for the Braun-Meise-Taylor setting and work within the general weight matrix setting. Finally, we construct a technical (counter-)example which is a weight in the sense of Beurling-Björck but violates the convexity condition and which illustrates the different flavor of both weight function settings.
Key words and phrases:
Weight functions, growth and regularity conditions for functions, test function spaces, convexity of functions
2020 Mathematics Subject Classification:
26A12, 26A48, 46A13, 46E10
This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/PAT9445424
1. Introduction
When dealing with ultradifferentiable function classes and analogously defined weighted spaces, in the literature “classically” one can find two settings: either using a weight sequence $\mathbf{M}$ or a weight function $\omega$. The ultradifferentiable weight sequence case is more original, we refer to [15] and [16]. On the other hand, in the weight function setting first the growth of the Fourier transform (of compactly supported functions) has been measured; see [1], [3], [23] and Section 6.3 for a summary. Here the expression measuring the admissible growth is given by $\omega$ directly (more precisely by $\exp\circ\omega$) and, indeed, non-trivial weighted test function spaces are treated (“Beurling-Björck setting”). However, in these original frameworks slightly different notions for being a weight function have been considered; we refer to Sections 2.8 and 6.3.
Later, in [7] the nowadays frequently used definition involving the Legendre-Fenchel-Young conjugate $\varphi^{*}_{\omega}$ of $\varphi_{\omega}:t\mapsto\omega(e^{t})$ has been given (“Braun-Meise-Taylor setting”). For such spaces the derivatives of the functions under investigation are weighted analogously to the weight sequence framework; see [7, Lemma 3.3, Prop. 3.4].
In each approach, basic growth and regularity assumptions on the weights $\omega$ are required and one can distinguish between two types: The Roumieu-type, involving the existence quantifier, and the Beurling-type when using the universal quantifier. As it has been mentioned in the introduction in [7], when comparing the assumptions between both settings it follows that requirement subadditivity for $\omega$ in [1], [3] and [23] can be replaced resp. relaxed by $\omega(2t)=O(\omega(t))$. This condition is frequently abbreviated by $(\alpha)$ and we write $(\omega_{1})$ for it; this abbreviation is motivated by the notation used in [29]. A further achievement is that the definition of weighted spaces within the Braun-Meise-Taylor setting admits the possibility to introduce and investigate the notion of quasianalyticity as well analogously to the weight sequence setting (“Denjoy-Carleman-Theorem”).
On the other hand, for establishing this equivalent description in [7, Lemma 3.3, Prop. 3.4] crucially it has been used that $\varphi_{\omega}$ is convex; usually abbreviated by $(\delta)$ and we write $(\omega_{4})$. This has become a standard assumption on $\omega$ also in related weighted settings since then; e.g. when considering globally defined functions of Gelfand-Shilov-type. $(\omega_{4})$ is crucial for obtaining the coincidence of both settings since the proof of the crucial Paley-Wiener-type theorem involves the double conjugate $\varphi^{**}_{\omega}$ and the identity $\varphi^{**}_{\omega}=\varphi_{\omega}$ which holds if and only if $\varphi_{\omega}$ is convex. In general, one only has $\varphi^{**}_{\omega}\leq\varphi_{\omega}$ and for this note that $\varphi^{**}_{\omega}$ and $\varphi^{*}_{\omega}$ both are automatically convex; see the comments in [29, Sect. 3.1] and more details are given in Section 2.9.
In [29] and [24] we have shown that ultradifferentiable classes in the Braun-Meise-Taylor setting can equivalently be described by using the associated weight matrix $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}:\ell>0\}$, $W^{(\ell)}_{j}:=\exp(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j))$, $j\in\mathbb{N}$; we refer to the main statement [24, Thm. 5.14] and for this alternative description $(\omega_{1})$ is crucial. It is immediate by definition that properties for $\omega$ can be transferred to $\mathcal{M}_{\omega}$ but, however, for the converse and hence when establishing a characterization of growth conditions, then again $(\omega_{4})$ is becoming relevant in order to treat the double conjugate. Another crucial property for which $(\omega_{4})$ is unavoidable is the equivalence between $\omega$ and each associated weight function $\omega_{\mathbf{W}^{(\ell)}}$; see [24, Lemma 5.7], [29, Theorem 4.0.3, Lemma 5.1.3], and also [13, Lemma 2.5]. On the other hand, for the sake of completeness we mention that basically $\mathcal{M}_{\omega}$ can be introduced for more general $\omega$. Indeed, it suffices to assume that $\omega$ is non-decreasing and $\log(t)=o(\omega(t))$ as $t\rightarrow+\infty$ (abbreviated by $(\omega_{3})$ in the following) in order to ensure well-definedness of the Young conjugate and hence of the sequences $\mathbf{W}^{(\ell)}$; we refer to the comments in [29, Sect. 3.1], [26, Sect. 2.7, (a)(c)(a)-(c)], [28, Sect. 6 $(III)$] and the summary in Section 2.9. Via $(\omega_{1})$ also the equality between the weighted spaces holds when the classes are introduced directly (by definition) via weighting the derivatives in terms of $\varphi^{*}_{\omega}$. Indeed, for this equality estimate [24, $(5.10)$] is required and for its proof $(\omega_{4})$ is not needed. For future reference we call any non-decreasing function satisfying $(\omega_{3})$ matrix admissible.
The main aim of this article is to illustrate the difference between the Beurling-Björck and the Braun-Meise-Taylor setting in view of the convexity for $\varphi_{\omega}$. Without this assumption the main Paley-Wiener-type result from [7] and hence the equality between both settings is becoming unclear. Therefore, when treating the pure Beurling-Björck framework the specific role resp. the loss of $(\omega_{4})$ has to be investigated. This question has been motivated by the explanations above and it is based on the comments in [29, Sect. 3.1]. But also in the recent article [20] in the introduction this convexity condition, its possible failure and its role in proofs within the Beurling-Björck are discussed. ($(\omega_{4})$ is also denoted by $(\delta)$ in [20].)
More precisely, by taking into account the fact that $(\omega_{4})$ is in general not preserved under equivalence of weight functions, we can see the first main question:
  • $(Q1)$
    How restrictive is $(\omega_{4})$? Does for any weight function (in the sense of Beurling-Björck) an equivalent function exist which satisfies $(\omega_{4})$?
Concerning the notion of equivalence of weight functions and more relevant growth relations we refer to Section 2.3. Connected to $(Q1)$ is the following problem:
  • $(Q2)$
    Characterize the inclusion relation, and hence the equality, of the spaces in the Beurling-Björck setting (for both types) in terms of the defining weights. One should expect the same growth relation between weight functions as in the Braun-Meise-Taylor setting (expressed by an O-growth restriction, written $\preceq$) and thus equivalence of weight functions characterizes the equality (as l.c.v.s).
We treat and answer both questions in this work in detail: The last Section 7 is dedicated to $(Q1)$ and to the construction of a (counter-)example showing that there exist weights in the sense of Beurling-Björck such that none equivalent (weight) function satisfies $(\omega_{4})$. Indeed, we even verify that there exist “many” of such weight functions; i.e. uncountable infinitely many different equivalence classes. The geometric construction also involves recent knowledge about the growth index $\gamma(\omega)$ for weight functions $\omega$ introduced in [12]; see Section 2.7. An immediate consequence of this example is to see the limitation of the weight sequence setting in this approach; i.e. when considering associated weight functions $\omega_{\mathbf{M}}$ with $\mathbf{M}$ being a sequence satisfying basic growth and regularity properties. This is due to the fact that $\omega_{\mathbf{M}}$ satisfies automatically $(\omega_{4})$; see Remark 7.6 for more details. On the other hand $\omega_{\mathbf{M}}$ appears frequently in the Braun-Meise-Taylor setting as a (counter-)example and in the comparison results from [6]; this fact again illustrates the difference between both settings. Finally, the constructed $\omega$ in Section 7 is matrix admissible.
$(Q2)$ is studied in detail in Sections 3, 4, 5 and 6; for the main characterizing results in this context see Theorem 4.1 for the Beurling-type and Theorem 4.2 for the Roumieu-type. We treat a very general setting: First, we work with weight function matrices $\mathcal{W}:=\{\omega^{\ell}:\ell>0\}$ with matrix parameter $\ell$, see Definition 3.1. Second, in Sections 3, 4 we are concerned with weighted subspaces of $L^{p}(\mathbb{R}^{d})$, $1\leq p\leq\infty$, subject to $\mathcal{W}$. Indeed, the weights are given by $x\mapsto\exp(\omega^{\ell}(x))$ and so involving the exponential function. Section 5 summarizes the situation when we focus on two special but natural weight function matrices for which the crucial information can be expressed in terms of a single weight function $\omega$: matrices of “exponential-type” and of “dilatation-type”. In Section 6 we apply the information to related and analogous situations: First, in Section 6.1 we consider the setting when weighting a fixed subspace $\mathcal{A}^{p}\subseteq L^{p}(\mathbb{R}^{d})$. In Section 6.2 we focus in detail on the case when weighting the “Fourier image” of a given (fixed) function space $\mathcal{A}$; i.e. not the growth of $f$ directly is controlled but of its Fourier transform $\widehat{f}$. Of course, for this case it is crucial to ensure that the Fourier transform is well-defined on $\mathcal{A}$ and its image is non-trivial. Finally, in each instance we also treat the anisotropic setting; i.e. when the weight functions are not radially extended to $\mathbb{R}^{d}$.
The known Beurling-Björck setting corresponds now to the special weight matrix $\omega^{\ell}:=\ell\omega$, $\ell>0$, where $\omega$ is given and fixed (“exponential type”), see Section 5.1, and when weighting the Fourier image of fixed subspaces (of functions having compact support) in $L^{1}(\mathbb{R}^{d})$. We refer to Section 6.3 for more details.
Before starting with this approach, in the first Section 2 we list, summarize and compare in detail the numerous growth conditions and relations on weight functions in the different settings. One can expect that this content is also interesting for other different weighted settings since we exclusively deal with growth and regularity properties for functions. The same comment applies to the construction provided in Section 7.
2. Weight functions and their growth conditions
2.1. General notation
We write $\mathbb{N}:=\{0,1,2,\dots\}$, $\mathbb{N}_{>0}:=\{1,2,\dots\}$ and for $x\in\mathbb{R}^{d}$ the expression $|x|$ denotes the usual Euclidean norm on $\mathbb{R}^{d}$.
2.2. Relevant growth properties
We list (known) conditions for functions $\omega:[0,+\infty)\rightarrow[0,+\infty)$; the abbreviations are also appearing in [29]. Some of these properties appear exclusively for weights in the sense of Beurling-Björck (see [3]), some of them only for Braun-Meise-Taylor weights (see [7]) and some of them are crucial in both settings.
We call a function $\omega$ normalized if $\omega(t)=0$ for all $0\leq t\leq 1$. Moreover, we shall assume $\omega\neq 0$ except stated explicitly otherwise. Inspired by the notation used in [29] we consider:
  • $(\omega_{1})$
    $\omega(2t)=O(\omega(t))$ as $t\rightarrow+\infty$.
  • $(\omega_{2})$
    $\omega(t)=O(t)$ as $t\rightarrow+\infty$.
  • $(\omega_{3})$
    $\log(t)=o(\omega(t))$ as $t\rightarrow+\infty$.
  • $(\omega_{4})$
    $\varphi_{\omega}:t\mapsto\omega(e^{t})$ is a convex function (on $\mathbb{R}$).
  • $(\omega_{5})$
    $\omega(t)=o(t)$ as $t\rightarrow+\infty$.
  • $(\omega_{6})$
    $\exists\;H\geq 1\;\forall\;t\geq 0:\;2\omega(t)\leq\omega(Ht)+H$.
  • $(\omega_{\text{nq}})$
    $\int_{1}^{\infty}\frac{\omega(t)}{t^{2}}dt<+\infty.$
  • $(\omega_{\text{snq}})$
    $\exists\;C>0\;\forall\;y>0:\int_{1}^{\infty}\frac{\omega(yt)}{t^{2}}dt\leq C\omega(y)+C$.
The next condition is crucial for characterizing desired stability properties for Braun-Meise-Taylor classes (of both types), see [24], [25] and the citations therein:
  • $(\alpha_{0})$
    $\exists\;C\geq 1\;\exists\;t_{0}\geq 0\;\forall\;\lambda\geq 1\;\forall\;t\geq t_{0}:\;\;\;\omega(\lambda t)\leq C\lambda\omega(t).$
In [29, $(3.8.4)$] this requirement has been denoted by $(\omega_{1^{{}^{\prime}}})$ but frequently in the literature it is abbreviated by $(\alpha_{0})$. Finally, let us introduce the following two requirements appearing in the Beurling-Björck framework (see [3]):
Finally, the Legendre-Fenchel-Young conjugate (of $\varphi_{\omega}$) is defined as follows:
(2.1) $\varphi^{*}_{\omega}(x):=\sup_{y\in\mathbb{R}}\{xy-\varphi_{\omega}(y)\},\;x\in\mathbb{R}.$
This transform is crucial when defining weighted classes in the sense of Braun-Meise-Taylor and is also basic for introducing the associated weight matrix $\mathcal{M}_{\omega}:=\{\mathbf{W}^{(\ell)}=(W^{(\ell)}_{j})_{j\in\mathbb{N}}:\ell>0\}$, $W^{(\ell)}_{j}:=\exp\left(\frac{1}{\ell}\varphi^{*}_{\omega}(\ell j)\right)$. We investigate this transform in more detail in Section 2.9.
2.3. Relevant growth relations between weight functions
Let $\sigma,\tau:[0,+\infty)\rightarrow[0,+\infty)$ be arbitrary functions and introduce the following growth relations between $\sigma$ and $\tau$:
    Write $\sigma\leq\tau$ if $\sigma(t)\leq\tau(t)$ for all $t\in[0,+\infty)$.
[…]
    $\exists\;C\geq 1\;\forall\;t\geq 0:\;\;\;\tau(t)\leq C\sigma(t)+C,$
[…]
    $\forall\;\epsilon>0\;\exists\;C\geq 1\;\forall\;t\geq 0:\;\;\;\tau(t)\leq\epsilon\sigma(t)+C.$
[…]
Note:
  • $(i)$
    In the literature, frequently the crucial growth relations for weight functions read as follows: $\sigma\preceq\tau$ resp. $\sigma\vartriangleleft\tau$ precisely means
    (2.2) $\tau(t)=O(\sigma(t))\;\text{as}\;t\rightarrow+\infty,\;\;\;\text{resp.}\;\;\;\tau(t)=o(\sigma(t))\;\text{as}\;t\rightarrow+\infty.$
    For these, in general stronger requirements, we also refer e.g. to [12, Rem. 2.6, $(5)$].
  • $(ii)$
    If $\sigma$ satisfies $\liminf_{t\rightarrow+\infty}\sigma(t)>0$; i.e.
    (2.3) $\exists\;t_{0}>0\;\exists\;\epsilon>0\;\forall\;t\geq t_{0}:\;\;\;\sigma(t)\geq\epsilon,$
[…]
  • $(iii)$
    (2.3) is a mild requirement; e.g. it holds if $\sigma\neq 0$ is assumed to be non-decreasing. Moreover it holds if $\lim_{t\rightarrow+\infty}\sigma(t)=+\infty$ and this has been a basic and natural growth requirement when studying weight functions even in a very general setting in [12].
  • $(iv)$
    If $\sigma:[0,+\infty)\rightarrow(0,+\infty)$, so if $\sigma(t)>0$ for all $t\geq 0$, then $\tau(t)=O(\sigma(t))$ as $t\rightarrow+\infty$ precisely means $\tau(t)\leq C\sigma(t)$ for some $C\geq 1$ and all $t\geq 0$; see e.g. the weighted entire setting treated in [31], [32].
All properties listed in Section 2.2 except $(\omega_{\operatorname{sub}})$ and the convexity property $(\omega_{4})$ are automatically preserved under $\sim$; see also [29, Sect. 3.2]. Indeed, the effects on $(\omega_{\operatorname{sub}})$ when passing to an equivalent weight and the relation to $(\alpha_{0})$ are investigated and revisited in Section 2.5 whereas the failure of $(\omega_{4})$ is studied in detail in Section 7 where a (counter-)example is constructed.
Naturally, one can also consider the following growth relations:
[…]
    $\exists\;C_{1},C_{2}\geq 1\;\forall\;t\geq 0:\;\;\;\tau(t)\leq\sigma(C_{1}t)+C_{2},$
[…]
    $\forall\;\epsilon>0\;\exists\;C\geq 1\;\forall\;t\geq 0:\;\;\;\tau(t)\leq\sigma(\epsilon t)+C.$
[…]
(2.4) $\forall\;u>0:\;\;\;\lim_{t\rightarrow+\infty}\frac{\sigma(tu)}{\sigma(t)}=1;$
[…]
2.4. On the comparison of crucial growth properties
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be given. We gather now immediate observations for the conditions listed in Section 2.2 and study relations between them.
  • $(a)$
    $(\omega_{3})$ is equivalent to $\lim_{t\rightarrow+\infty}\frac{\varphi_{\omega}(t)}{t}=+\infty$ and it clearly implies $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$. Moreover, $(\omega_{3})$ is equivalent to $\log(a+t)=o(\omega(t))$ as $t\rightarrow+\infty$ and $(\omega_{3,w})$ is equivalent to $\log(a+t)=O(\omega(t))$ as $t\rightarrow+\infty$, where $a\geq 1$ is an arbitrary real parameter.
  • $(b)$
    $(\omega_{3})$ implies $(\omega_{3,w})$ and $(\omega_{3,w})$ implies $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$. The crucial difference between these two conditions is that $(\omega_{3,w})$ formally allows to consider the “limiting weight” $\omega(t):=\log(1+t)$ whereas $(\omega_{3})$ excludes this special case. In the literature occasionally $t\mapsto\log(1+t)$ is treated for the Beurling-type and which gives back indeed the non-weighted situation.
  • $(c)$
  • $(d)$
    $(\alpha_{0})$ implies $(\omega_{1})$ (choose $\lambda=2$).
  • $(e)$
    Also $(\omega_{\operatorname{sub}})$ implies $(\omega_{1})$ (take $x=y$) and with $x=y=0$ we get $\omega(0)\leq 2\omega(0)$ which yields $\omega(0)\geq 0$.
  • $(f)$
    In the literature sometimes the following condition is used (see e.g. [7, Lemma 1.2]):
    (2.5) $\exists\;L\geq 1\;\forall\;s,t\geq 0:\;\;\;\omega(s+t)\leq L(\omega(s)+\omega(t))+L.$
    Set $s=t$ and so get immediately $(\omega_{1})$ when choosing the constant $2L$. If $\omega$ is non-decreasing, then (2.5) is equivalent to $(\omega_{1})$: Indeed, if $\omega$ has $(\omega_{1})$ then $\omega(s+t)\leq\omega(2\max\{s,t\})\leq L\omega(\max\{s,t\})+L\leq L(\omega(s)+\omega(t))+L$ and so (2.5) holds with the same $L$.
  • $(g)$
  • $(h)$
    If $\omega$ is non-decreasing, then $(\omega_{\operatorname{nq}})$ implies $(\omega_{5})$ because $\int_{s}^{\infty}\frac{\omega(t)}{t^{2}}dt\geq\omega(s)\int_{s}^{\infty}\frac{1}{t^{2}}dt=\frac{\omega(s)}{s}$ for all $s\geq 1$.
  • $(i)$
    Let $\omega$ be non-decreasing (and such that $\omega\neq 0$). Then concavity on $0,+\infty)$ and normalization can never be satisfied at the same time.
    The same technical problem appears for $(\alpha_{0})$, see also [30, Sect. 4.1] and the references therein: Assume that there exists $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$, then $C\lambda\omega(t)=0$ for all $C,\lambda\geq 1$ and all $0<t\leq t_{0}$. But, since $\omega\neq 0$ and $\omega$ is non-decreasing, we have $\omega(\lambda t)>0$ as $\lambda\rightarrow+\infty$ for any $t>0$.
    And similarly also normalization and $(\omega_{\operatorname{sub}})$ can never be achieved at the same time: If there exists $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$ then for all $0<s\leq t_{0}$ and all $t\geq 0$ we get $\omega(t)\leq\omega(s+t)\leq\omega(s)+\omega(t)=\omega(t)$. But this implies $\omega=0$.
We close this section by investigating conditions $(\omega_{1})$ and $(\omega_{6})$. First, we state a technical result.
Lemma 2.1.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing.
  • $(I)$
    The following are equivalent:
    • $(i)$
      $\omega$ satisfies $(\omega_{1})$.
    • $(ii)$
      $\omega$ satisfies
      $\exists\;a>1\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\omega(at)\leq L\omega(t)+L.$
    • $(iii)$
      $\omega$ satisfies
      $\forall\;a>1\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\omega(at)\leq L\omega(t)+L$
  • $(II)$
    The following are equivalent:
    • $(i)$
      $\omega$ satisfies $(\omega_{6})$.
    • $(ii)$
      $\omega$ satisfies
      $\exists\;a>1\;\exists\;H\geq 1\;\forall\;t\geq 0:\;\;\;a\omega(t)\leq\omega(Ht)+H.$
    • $(iii)$
      $\omega$ satisfies
      $\forall\;a>1\;\exists\;H\geq 1\;\forall\;t\geq 0:\;\;\;a\omega(t)\leq\omega(Ht)+H.$
Note that assumption [10, $(\alpha)$] precisely corresponds to $(I)(ii)$.
Proof. The particular implications $(i)\Rightarrow(ii)$ and $(iii)\Rightarrow(i)$ are clear (set $a:=2$) and $(ii)\Rightarrow(iii)$ follows by iteration; concerning $(\omega_{6})$ see also [28, Rem. 2.2]. However, the assumption $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ is not required necessarily for these equivalences. ∎
[…]
Proposition 2.2.
Let $\sigma,\tau:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing.
  • $(i)$
[…]
  • $(ii)$
[…]
  • $(iii)$
[…]
  • $(iv)$
[…]
2.5. On subadditivity-like conditions
We recall the following equivalent reformulations of $(\alpha_{0})$; see [21, Lemma 1], [12, Prop. 2.23] and finally [30, Lemma 4.1]:
Lemma 2.3.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing. Then the following are equivalent:
  • $(i)$
    $\omega$ is equivalent to a subadditive function $\sigma$.
  • $(ii)$
    $\omega$ satisfies $(\alpha_{0})$.
  • $(iii)$
    $\omega$ satisfies
    (2.6) $\exists\;C\geq 1\;\exists\;D\geq 1\;\forall\;\lambda\geq 1\;\forall\;t\geq 0:\;\;\;\omega(\lambda t)\leq C\lambda\omega(t)+D\lambda.$
  • $(iv)$
    $\omega$ is equivalent to its least concave majorant $F_{\omega}$.
Note:
    Assumption $\omega(0)\geq 0$ in [30, Lemma 4.1] is superfluous.
    To be formally correct, in the proof of [30, Lemma 4.1 $(ii)\Rightarrow(iii)$] one shall assume $s>0$ and note that the second part of [12, $(12)$] is trivial when $s=0$.
    In [30] equivalence is defined via O-growth relations like in (2.2), see [30, Sect. 2.4, p. 406], and which coincides with the above given notion $\sim$ since (2.3) holds by $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ as a basic requirement in this work.
    However, in [30, Lemma 4.1] it was only assumed that $\omega$ is non-decreasing but this is sufficient to get (2.3) when $\omega\neq 0$ and as seen before both notions of equivalence coincide in this case.
    And if $\omega=0$, then one immediately infers $F_{\omega}=0$ too and the above result is trivial.
$(\alpha_{0})$ should be compared with $(\alpha_{1})$ in [23]. Note that in [23, Prop. 1.1] it has been claimed that $(\alpha_{1})$ is equivalent to $(\alpha_{0})$ but which fails in general and $(\alpha_{1})$ is weaker. However, the difference between both requirements is subtle, we refer to [30] for more comments and explanations.
In the next step we show that (2.6) is also related to the largest convex minorant; see the proofs of [12, Prop. 2.24 & Cor. 2.26].
For arbitrary $\omega:[0,+\infty)\rightarrow[0,+\infty)$ introduce $\omega^{\iota}:(0+\infty)\rightarrow[0,+\infty)$ given by $\omega^{\iota}(t):=\omega(1/t)$. So $\omega^{\iota}$ is non-increasing if and only if $\omega$ is non-decreasing, $\lim_{t\rightarrow 0}\omega^{\iota}(t)=\lim_{t\rightarrow+\infty}\omega(t)$ and $\lim_{t\rightarrow+\infty}\omega^{\iota}(t)=\lim_{t\rightarrow 0}\omega(t)$ if the limits exist.
$\overline{\omega^{\iota}}^{c}$ shall denote the largest convex minorant of $\omega^{\iota}$ which can be represented as follows, see e.g. the proof of [12, Prop. 2.24]:
$\overline{\omega^{\iota}}^{c}(t):=\inf\{\lambda_{1}\omega^{\iota}(t_{1})+\lambda_{2}\omega^{\iota}(t_{2}):\;\lambda_{1}+\lambda_{2}=1,\;\lambda_{1}t_{1}+\lambda_{2}t_{2}=t,\;\lambda_{i}\geq 0\}.$
Of course, this definition makes sense for any function $h:(0+\infty)\rightarrow[0,+\infty)$ and $\overline{h}^{c}(t)\leq h(t)$ for all $t$ when choosing $\lambda_{1}=1$, $\lambda_{2}=0$ and $t_{1}=t$, $t_{2}\in(0,+\infty)$ arbitrary.
Proposition 2.4.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing. If $\omega$ satisfies (2.6), then
$\exists\;A\geq 1\;\forall\;t\in(0,+\infty):\;\;\;\overline{\omega^{\iota}}^{c}(t)\leq\omega^{\iota}(t)\leq A\overline{\omega^{\iota}}^{c}(t)+A;$
hence $\omega^{\iota}$ is equivalent to its largest convex minorant.
Proof. We verify [12, $(14)$] for $\omega^{\iota}$, i.e.
(2.7) $\exists\;C\geq 1\;\exists\;\beta>0\;\forall\;s,t\in(0,+\infty):\;\;\;\omega^{\iota}(s)+C\geq\frac{1}{C}\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(t),$
and then apply [12, Prop. 2.24]. Note that (2.7) is equivalent to the existence of universal constants $C_{1},C,D\geq 1$ not depending on $\beta$ such that $C_{1}\omega^{\iota}(s)+D\geq\frac{1}{C_{2}}\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(s)$ holds for all $s,t\in(0,+\infty)$ (set $C:=\max\{D/C_{1},C_{1}C_{2}\}$).
First, since $\omega$ is non-decreasing for any $0<s\leq t$ we have $\omega^{\iota}(t)=\omega(1/t)\leq\omega(1/s)=\omega^{\iota}(s)$ and $\min\{1,\frac{t^{\beta}}{s^{\beta}}\}=1$ for all $\beta>0$. Thus (2.7) holds in this situation with any $C\geq 1$ and $\beta>0$.
Second, note that (2.6) gives with $t^{\prime}:=\frac{1}{\lambda t}$:
$\exists\;C,D\geq 1\;\forall\;\lambda\geq 1\;\forall\;t^{\prime}>0:\;\;\;\omega^{\iota}(t^{\prime})=\omega(1/t^{\prime})\leq C\lambda\omega(1/(\lambda t^{\prime}))+D\lambda=C\lambda\omega^{\iota}(\lambda t^{\prime})+D\lambda.$
Let now $0<t<s$ be arbitrary and choose $\lambda=\frac{s}{t}>1$, then the above estimate gives
$D\lambda+C\lambda\omega^{\iota}(s)=D\lambda+C\lambda\omega^{\iota}(\lambda t)\geq\omega^{\iota}(t),$
hence for any $\beta\geq 1$:
$D+C\omega^{\iota}(s)\geq\frac{t}{s}\omega^{\iota}(t)\geq\frac{t^{\beta}}{s^{\beta}}\omega^{\iota}(t)=\min\left\{1,\frac{t^{\beta}}{s^{\beta}}\right\}\omega^{\iota}(t).$
Consequently, (2.7) is shown and [12, Prop. 2.24] yields the assertion. ∎
On the other hand, if there exists a convex function $\sigma:(0,+\infty)\rightarrow[0,+\infty)$ which is equivalent to $\omega^{\iota}$, then there exists some $A\geq 1$ such that for all $t,s\in(0,+\infty)$ and $\lambda_{1},\lambda_{2}\in[0,1]$ satisfying $\lambda_{1}+\lambda_{2}=1$:
$\omega^{\iota}(\lambda_{1}t+\lambda_{2}s)$ $\leq A\sigma(\lambda_{1}t+\lambda_{2}s)+A\leq A\lambda_{1}\sigma(t)+A\lambda_{2}\sigma(s)+A$
$\leq A^{2}\lambda_{1}\omega^{\iota}(t)+A^{2}\lambda_{1}+A^{2}\lambda_{2}\omega^{\iota}(s)+A^{2}\lambda_{2}=A^{2}\lambda_{1}\omega^{\iota}(t)+A^{2}\lambda_{2}\omega^{\iota}(s)+A^{2}.$
And if there exists some $t_{0}>0$ such that $\omega(t)=0$ for all $t\in[0,t_{0}]$, then $\omega^{\iota}(t)=0$ for all $t\geq t_{0}^{-1}$. Applying the previous estimate to $s:=t_{0}^{-1}$ gives $\omega^{\iota}(\lambda_{1}t+\lambda_{2}t_{0})\leq A^{2}\lambda_{1}\omega^{\iota}(t)+A^{2}$ for all $t\in(0,t_{0}]$ and $\lambda_{1},\lambda_{2}\geq 0$ satisfying $\lambda_{1}+\lambda_{2}=1$.
Remark that the proof and the conclusion from [29, Lemma 3.8.1 $(2)$] fails since super-additivity for $\omega^{\iota}$ would imply $2\omega^{\iota}(t)\leq\omega^{\iota}(2t)\leq\omega^{\iota}(t)$ for all $t\in(0,+\infty)$, a contradiction.
2.6. Comment on weights being defined on $\mathbb{R}^{d}$ or $\mathbb{C}^{d}$
Usually, when being defined on $\mathbb{R}^{d}$, $d\geq 2$, it is assumed that $\omega$ is extended radially and so $\omega(x):=\omega(|x|)$ for any $x\in\mathbb{R}^{d}$ (isotropic setting). On the other hand, for $z\in\mathbb{C}^{d}$, in the literature one can find the definition $\omega(z):=\omega(|z|_{1})$ with $|z|_{1}:=\sum_{j=1}^{d}|z_{j}|$ for $z=(z_{1},\dots,z_{d})$; see e.g. [7, Def. 1.1].
Write $z_{j}=x_{j}+iy_{j}$, $1\leq j\leq d$, then $|z|_{1}=\sum_{j=1}^{d}(x_{j}^{2}+y_{j}^{2})^{1/2}$ and $|z|=\left(\sum_{j=1}^{d}|z_{j}|^{2}\right)^{1/2}=\left(\sum_{j=1}^{d}x_{j}^{2}+y_{j}^{2}\right)^{1/2}$, so $|z|\leq|z|_{1}\leq d\max\{|z_{j}|:1\leq j\leq d\}\leq d|z|$ holds for all $z\in\mathbb{C}^{d}$.
Using these estimates we see that for all non-decreasing $\omega$ satisfying $(\omega_{1})$ both notions are equivalent in the sense that the mappings $z\mapsto\omega(|z|)$ and $z\mapsto\omega(|z|_{1})$ are equivalent w.r.t. relation $\sim$.
More recently also non-radial weights have been considered (anisotropic setting); we refer e.g. to [20] and [8].
2.7. The growth index $\gamma(\omega)$
We briefly recall the definition of the growth index $\gamma(\omega)$, see [12, Sect. 2.3] and the references therein. Let in this section $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing and such that $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$; i.e. $\omega$ is a weight function as in [12, Sect. 2.3]. For $\gamma>0$ we say that $\omega$ has property $(P_{\omega,\gamma})$ if
$\exists\;K>1:\;\;\;\limsup_{t\rightarrow+\infty}\frac{\omega(K^{\gamma}t)}{\omega(t)}<K.$
If $(P_{\omega,\gamma})$ holds for some $K>1$, then also $(P_{\omega,\gamma^{\prime}})$ is satisfied for all $\gamma^{\prime}\leq\gamma$ with the same $K$ since $\omega$ is non-decreasing. Moreover, we can restrict to $\gamma>0$, because for $\gamma\leq 0$ condition $(P_{\omega,\gamma})$ is automatically satisfied (again since $\omega$ is non-decreasing and $K>1$). Then we put
(2.8) $\gamma(\omega):=\sup\{\gamma>0:(P_{\omega,\gamma})\;\;\text{is satisfied}\},$
and if none condition $(P_{\omega,\gamma})$ holds then set $\gamma(\omega):=0$.
Remark 2.5.
Let $\omega$ be as before. We recall some connections between $\gamma(\omega)$ and sub-additivity-like conditions for $\omega$, see also [30, Sect. 6]:
    By [12, Cor. 2.14] we get $\gamma(\omega)>0$ if and only if $(\omega_{1})$ is valid.
    By [12, Thm. 2.11, Cor. 2.13] we have that $\gamma(\omega)>1$ if and only if $\omega$ has $(\omega_{\operatorname{snq}})$.
[…]
    (2.9) $\kappa_{\omega}(y):=\int_{1}^{+\infty}\frac{\omega(yt)}{t^{2}}dt.$
    $\kappa_{\omega}$ is concave, continuous and we have $\kappa_{\omega}(0)=\omega(0)\geq 0$ (i.e. $\kappa_{\omega}(0)=0$ if and only if $\omega(0)=0$). Hence it is known that also $(\omega_{\operatorname{sub}})$ holds for $\kappa_{\omega}$; see e.g. [29, Lemma 3.8.1 $(1)$].
    Note: If continuity for $\omega$ fails, then the proof of [12, Thm. 2.11 $(i)\Rightarrow(ii)$] yields the fact that we get all desired properties for $y\mapsto\widetilde{\kappa}_{\omega}(y):=\int_{1}^{+\infty}\frac{\kappa_{\omega}(yt)}{t^{2}}dt$ and so $\omega$ in (2.9) is replaced by $\kappa_{\omega}$. Consequently, in any case one has that $\omega$ is equivalent to a continuous weight satisfying $(\omega_{\operatorname{sub}})$.
The comments listed in Section 2.4, Remark 2.5 and Lemma 2.3 give the following result:
Lemma 2.6.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing and such that $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$. Then the following implications hold
[…]
and, moreover,
[…]
2.8. Different notions of weight functions
In this section we revisit the different notions of weight functions appearing in the literature.
Definition 2.7.
$\omega:[0,+\infty)\rightarrow[0,+\infty)$ is called a weight function in the terminology of [12, Sect. 2.1], [13, Sect. 2.2] and [14, Sect. 2.2], if $\omega$ is
    continuous,
    non-decreasing,
    $\omega(0)=0$, and
    $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$.
If $\omega$ satisfies in addition $\omega(t)=0$ for all $t\in[0,1]$, then we call $\omega$ normalized.
Note: Occasionally in [12] we have even worked with a more general notion. Indeed, the assumptions continuity and $\omega(0)=0$ can be removed when working with abstract growth indices; see [12, Sect. 2.2-2.5] and Section 2.7.
Definition 2.8.
$\omega:[0,+\infty)\rightarrow[0,+\infty)$ is called a weight function in the sense of Braun-Meise-Taylor, see [7, Def. 1.1] and write BMT-weight for short, if
Moreover:
    If $\omega$ satisfies in addition $\omega(t)=0$ for all $t\in[0,1]$, then $\omega$ is called normalized.
    If $\omega$ has in addition $(\omega_{\operatorname{nq}})$, then $\omega$ is called non-quasianalytic and quasianalytic otherwise.
    If $(\omega_{\operatorname{snq}})$ holds then $\omega$ is called strong non-quasianalytic or even strong for short. This terminology is inspired by [5, Def. 1.8].
Remark 2.9.
We gather some comments on the assumptions of BMT-weights:
    In [7, Def. 1.1] property $(\omega_{\operatorname{nq}})$ and normalization have been standard assumptions.
    It is also known that normalization can be assumed w.l.o.g. since the corresponding weighted classes are preserved under equivalence; for the ultradifferentiable classes $\mathcal{E}_{\{\omega\}}$ resp. $\mathcal{E}_{(\omega)}$ see [5, Rem. 1.2 $(b)$] and [24, Cor. 5.17].
    Formally, the basic requirement $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$ can be skipped since it follows in any case by $(\omega_{3})$.
    Based on the notation [7, Def. 1.1], in the literature when dealing with BMT-weight functions $(\omega_{1})$ is frequently denoted by $(\alpha)$, $(\omega_{\operatorname{nq}})$ by $(\beta)$, $(\omega_{3})$ by $(\gamma)$ and $(\omega_{4})$ by $(\delta)$.
    According to [18, Sect. 2.1] and [19, Sect. 2.2] a (normalized) BMT-weight $\omega$ is called a (normalized) weight function and when only $(\omega_{1})$ fails then it is called a (normalized) pre-weight function.
Next we move to the Beurling-Björk setting.
Definition 2.10.
According to [3, Def. 1.3.22] the function $\omega:[0,+\infty)\rightarrow[0,+\infty)$ is called a weight in the sense of Beurling-Björk, and write BB-weight for short, if $\omega$ is
More precisely, in [3]
    the set of functions from Definition 2.10 is denoted by $\mathfrak{M}$ and $\omega$ has been radially extended to $\mathbb{R}^{d}$;
    requirements $(ii)$ and $(iii)$ together are abbreviated by $(\alpha)$, $(iv)$ by $(\beta)$ and $(v)$ by $(\gamma)$;
    according to [3, Def. 1.2.5] the set $\mathfrak{M}_{0}$ denotes functions satisfying (i)(iv)(i)-(iv).
For the sake of completeness let us mention that also in [23] and [7, Sect. 8] the notion of BB-weights has been considered; in [7] for a comparison of different theories.
More precisely, in [7, 8.4 $(2)$] for weights in the sense of Beurling-Björck all requirements from Definition 2.10 except (formally) continuity have been considered.
Finally, in [23] H.-J. Petzsche and D. Vogt have assumed that
Occasionally, the authors have also considered
Note: $(\delta)$ in [23, Sect. 5] does not mean $(\omega_{4})$ (i.e. the convexity of $\varphi_{\omega}$) and $(\gamma_{1})$ must not be mixed and confused with Petzsche’s condition for weight sequences introduced in [22].
2.9. Matrix admissible weight functions
[…]
  • $(a)$
    If $\omega$ is normalized, then $\sup_{y\in\mathbb{R}}\{xy-\varphi_{\omega}(y)\}=\sup_{y\geq 0}\{xy-\varphi_{\omega}(y)\}$ for all $x\geq 0$. Note that in concrete applications, e.g. in the BMT-setting, it is natural to consider $x\in[0,+\infty)$ in (2.1) but, however, naturally one should consider $\sup_{y\in\mathbb{R}}$ in the definition since $\mathbb{R}$ is the natural domain of definition for $\varphi_{\omega}$.
  • $(b)$
    By definition $\varphi^{*}_{\omega}$ is convex and $\varphi^{*}_{\omega}(0)=\sup_{y\in\mathbb{R}}\{-\varphi_{\omega}(y)\}$ and so, if $\inf_{t\geq 0}\omega(t)=0$, then $\varphi^{*}_{\omega}(0)=0$ holds and if $\omega$ is non-decreasing, then $\varphi^{*}_{\omega}(0)=-\omega(0)$.
    Thus, if $\omega$ is normalized, then $\varphi^{*}_{\omega}(0)=0=\omega(0)$ and, moreover, $\varphi^{*}_{\omega}$ is non-decreasing on $0,+\infty)$ (see $(a)$). But, however, in general this property is not satisfied.
  • $(c)$
    Naturally, one should restrict $\varphi^{*}_{\omega}$ to $0,+\infty)$: Let $x<0$, then
    $\varphi^{*}_{\omega}(x)=\sup_{y\in\mathbb{R}}\{xy-\varphi_{\omega}(y)\}\geq\sup_{y\in(-\infty,0]}\{xy-\varphi_{\omega}(y)\}\geq\sup_{y\in(-\infty,0]}\{xy\}-\sup_{y\in(-\infty,0]}\{\varphi_{\omega}(y)\}=+\infty.$
    For this note that $\omega(t)\in[0,+\infty)$ for all $t\in[0,1]$ and so $\varphi^{*}_{\omega}=+\infty$ on $(-\infty,0)$.
  • $(d)$
    $x\mapsto\frac{\varphi^{*}_{\omega}(x)}{x}$ is non-decreasing on $(0,+\infty)$ and $\lim_{x\rightarrow+\infty}\frac{\varphi^{*}_{\omega}(x)}{x}=+\infty$: The first property follows because $y-\frac{\varphi_{\omega}(y)}{x_{1}}\leq y-\frac{\varphi_{\omega}(y)}{x_{2}}\Leftrightarrow\frac{1}{x_{2}}\leq\frac{1}{x_{1}}$ for all $0<x_{1}\leq x_{2}$ and $y\in\mathbb{R}$. Concerning the second one assume that $\frac{\varphi^{*}_{\omega}(x)}{x}\leq C$ for some $C>0$ and all $x>0$, then by definition $x(y-C)\leq\varphi_{\omega}(y)$ for all $x\geq 0$ and $y\in\mathbb{R}$. But this is impossible: Choose a fixed $y>C$ and then let $x\rightarrow+\infty$.
  • $(e)$
    $(\omega_{3})$ for $\omega$ precisely means $\lim_{y\rightarrow+\infty}\frac{\varphi_{\omega}(y)}{y}=+\infty$ and this property characterizes $\varphi^{*}_{\omega}(x)<+\infty$ for all $x\in[0,+\infty)$: Indeed, if $(\omega_{3})$ is valid then for any $C>0$ (large) we can find $d\in\mathbb{R}$, w.l.o.g. $d<0$, such that $\varphi_{\omega}(y)\geq yC+d\Leftrightarrow-d\geq yC-\varphi_{\omega}(y)$ for all $y\geq 0$. Since $\varphi_{\omega}(y)\geq 0$ for any $y\in\mathbb{R}$ and because $C>0$ the same estimate holds for any $y\in\mathbb{R}$. Hence $\varphi^{*}_{\omega}(C)\leq-d$ follows. Since $C>0$ is arbitrary $\varphi^{*}_{\omega}$ is well-defined on $0,+\infty)$. Conversely, if $\varphi^{*}_{\omega}$ is well-defined on $0,+\infty)$, then for all $x\geq 0$ there exists $C_{x}>0$ such that $xy-\varphi_{\omega}(y)\leq C_{x}$ for all $y\in\mathbb{R}$ and so $x-\frac{C_{x}}{y}\leq\frac{\varphi_{\omega}(y)}{y}$ for all $y>0$. When $x\geq 0$ is fixed, then this estimate verifies $\liminf_{y\rightarrow+\infty}\frac{\varphi_{\omega}(y)}{y}\geq x$ and as $x\rightarrow+\infty$ we infer $(\omega_{3})$ for $\omega$.
  • $(f)$
    Now consider the double conjugate $\varphi^{**}_{\omega}$: This function is by definition also convex and, by $(c)$, one has $\varphi^{**}_{\omega}(x)=\sup_{y\in\mathbb{R}}\{xy-\varphi^{*}_{\omega}(y)\}=\sup_{y\geq 0}\{xy-\varphi^{*}_{\omega}(y)\}$ for all $x\in\mathbb{R}$. Comments $(d)$ and $(e)$ together imply that $\varphi^{**}_{\omega}$ is always well-defined (on $\mathbb{R}$): Indeed, $(d)$ gives $(\omega_{3})$ for $\omega^{*}:=\varphi^{*}_{\omega}\circ\log$ and so well-definedness on $0,+\infty)$ whereas on $(-\infty,0)$ this follows since $\varphi^{**}_{\omega}$ is non-decreasing.
    This fact does also follow from $\varphi^{**}_{\omega}(x)=\sup_{y\in\mathbb{R}}\{xy-\varphi^{*}_{\omega}(y)\}=\sup_{y\in\mathbb{R}}\{xy-\sup_{z\in\mathbb{R}}\{zy-\varphi_{\omega}(z)\}\}\leq\varphi_{\omega}(x)$, $x\in\mathbb{R}$ arbitrary.
    Concerning the converse, note that $\varphi^{**}_{\omega}=\varphi_{\omega}$ clearly implies $(\omega_{4})$ for $\omega$. If $\omega$ is in addition non-decreasing, then we show that $(\omega_{4})$ implies $\varphi^{**}_{\omega}\geq\varphi_{\omega}$ and so $\varphi^{**}_{\omega}=\varphi_{\omega}$. The following arguments are inspired by (the proof of) [23, Prop. 1.6]:
    Let $t\in\mathbb{R}$ be fixed and since $\varphi_{\omega}$ is convex one has that $u\mapsto\frac{\varphi_{\omega}(u)-\varphi_{\omega}(t)}{u-t}$ is non-decreasing, $u\in\mathbb{R}$ such that $u\neq t$. Since $\varphi_{\omega}$ is non-decreasing this difference quotient is non-negative. So there exists $s=s(t)\geq 0$ such that
    $\inf_{u>t}\frac{\varphi_{\omega}(u)-\varphi_{\omega}(t)}{u-t}=(\varphi_{\omega})^{\prime}_{+}(t)\geq s\geq(\varphi_{\omega})^{\prime}_{-}(t)=\sup_{u<t}\frac{\varphi_{\omega}(u)-\varphi_{\omega}(t)}{u-t}.$
    Multiplying these estimates with utu-t we get that $s(u-t)\leq\varphi_{\omega}(u)-\varphi_{\omega}(t)$ for all $u\in\mathbb{R}$ and note that for $u=t$ this estimate is becoming trivial. Consequently, $\inf_{s\in\mathbb{R}}\sup_{u\in\mathbb{R}}\{s(u-t)-\varphi_{\omega}(u)\}\leq-\varphi_{\omega}(t)$ holds which is equivalent to having $\inf_{s\in\mathbb{R}}\sup_{u\in\mathbb{R}}\{s(u-t)-\varphi_{\omega}(u)\}\geq\varphi_{\omega}(t)$ for all $t\in\mathbb{R}$. The left-hand side is equal to $\sup_{s\in\mathbb{R}}\{-\sup_{u\in\mathbb{R}}\{s(u-t)-\varphi_{\omega}(u)\}\}=\sup_{s\in\mathbb{R}}\{st-\sup_{u\in\mathbb{R}}\{su-\varphi_{\omega}(u)\}\}=\varphi_{\omega}^{**}(t)$ and we are done.
    Note: If $\omega$ is normalized, non-decreasing and satisfies $(\omega_{4})$, then by the above and $(b)$ we get $(0=)-\varphi^{*}_{\omega}(0)=\omega(0)=\omega(1)=\varphi_{\omega}(0)=\varphi^{**}_{\omega}(0)=\sup_{y\in\mathbb{R}}\{-\varphi^{*}_{\omega}(y)\}=\sup_{y\geq 0}\{-\varphi^{*}_{\omega}(y)\}$ and for having $(0=)-\varphi^{*}_{\omega}(0)=\sup_{y\geq 0}\{-\varphi^{*}_{\omega}(y)\}$ it suffices to assume normalization since then $\varphi^{*}_{\omega}$ is non-decreasing on $0,+\infty)$; see again $(a)$ and $(b)$.
Summarizing, assume that $\omega$ is non-decreasing and such that $(\omega_{3})$ holds. Then $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$, so $\omega$ is a weight function in the notion of [12] (and [27], [26]), and by (a)(e)(a)-(e) it follows that each $\mathbf{W}^{(\ell)}$ is a well-defined sequence of non-negative real numbers, each $\mathbf{W}^{(\ell)}$ is log-convex (i.e. $(M.1)$ in [15] is valid), $\lim_{p\rightarrow+\infty}(W^{(\ell)}_{p})^{1/p}=+\infty$ for all $\ell>0$, and finally $W^{(\ell_{1})}_{p}\leq W^{(\ell_{2})}_{p}$ for all $p\in\mathbb{N}$, $0<\ell_{1}\leq\ell_{2}$. Thus basically “non-decreasing” and $(\omega_{3})$ for $\omega$ are sufficient to introduce $\mathcal{M}_{\omega}$ and this explains the notion “matrix admissible”.
If $\omega$ is also normalized, then each $\mathbf{W}^{(\ell)}$ is normalized; i.e. $1=W^{(\ell)}_{0}\leq W^{(\ell)}_{1}$. Note that, when $\omega$ is not normalized, then one can always switch to an equivalent weight $\sigma$ which is normalized and this notion of equivalence transfers in a precise manner to the corresponding associated weight matrices and which ensures the equality of the corresponding (matrix) weighted spaces; we refer to [28, Sect. 6 $(III)$, $(6.6)$].
Finally, $(\omega_{4})$ gives the equality $\varphi^{**}_{\omega}=\varphi_{\omega}$ and ensures the equivalence between $\omega$ and (each) $\omega_{\mathbf{W}^{(\ell)}}$, see [24, Lemma 5.7]. When $\omega$ is matrix admissible but not having $(\omega_{4})$ then this equivalence is not true in general: This follows by the example constructed in Section 7 and recall that each associated weight function satisfies $(\omega_{4})$, see also Remark 7.6.
2.10. The (counter-)example by U. Franken
In [10, Prop. 3] the following statement has been shown:
Theorem 2.11.
There exists a non-quasianalytic BMT-weight function $\omega$ such that each continuous increasing function $\sigma$ with $\sigma\geq\omega$ cannot have simultaneously $(\omega_{\operatorname{sub}})$ and $(\omega_{\operatorname{nq}})$ and hence cannot be a weight in the sense of Beurling-Björck.
3. Weighted $L^{p}$-type spaces
3.1. Basic definitions
For functions $\sigma,\tau:\mathbb{R}^{d}\rightarrow[0,+\infty)$ we write $\sigma\leq\tau$ if $\sigma(x)\leq\tau(x)$ for all $x\in\mathbb{R}^{d}$ and similarly all further relations from Section 2.3 transfer. We give the following general definition:
Definition 3.1.
Consider the set of functions
$\mathcal{W}:=\{\omega^{\ell}:\mathbb{R}^{d}\rightarrow[0,+\infty),\;\;\;\ell>0:\;\;\;\omega^{\ell_{1}}\leq\omega^{\ell_{2}},\;\forall\;0<\ell_{1}\leq\ell_{2}\}.$
$\mathcal{W}$ is denoted as a weight function matrix and $\ell>0$ is the crucial matrix index parameter. $\mathcal{W}$ is called
    simple if $\mathcal{W}=\{\omega\}$ which means that $\omega^{\ell}\equiv\omega$ for any $\ell>0$;
    radial or isotropic if $\omega^{\ell}(x)=\omega^{\ell}(|x|)$ for $x\in\mathbb{R}^{d}$ and $\ell>0$;
    non-decreasing if all $\omega^{\ell}$ are non-decreasing which means $\omega^{\ell}(x)\leq\omega^{\ell}(y)$ for all $x,y\in\mathbb{R}^{d}$ such that $|x|\leq|y|$;
    continuous if all $\omega^{\ell}$ are continuous,
    unbounded if
    (3.1) $\forall\;\ell>0:\;\;\;\sup_{x\in\mathbb{R}^{d}}\omega^{\ell}(x)=+\infty.$
If $\mathcal{W}$ is non-decreasing, then $\sup_{x\in\mathbb{R}^{d}:|x|=t}\omega^{\ell}(x)\leq\inf_{\epsilon>0}\inf_{x\in\mathbb{R}^{d}:|x|=t+\epsilon}\omega^{\ell}(x)$ for any $t\geq 0$ fixed (and any $\ell>0$). Therefore, if $\mathcal{W}$ is in addition continuous, then $\mathcal{W}$ has to be radial; i.e. each $\omega^{\ell}$ is constant on each sphere $|x|=t$.
If $\mathcal{W}$ is radial, non-decreasing and unbounded, then $\lim_{|x|\rightarrow+\infty}\omega^{\ell}(x)=+\infty$ for all $\ell>0$.
For any $1\leq p\leq\infty$ set $L^{p}=L^{p}(\mathbb{R}^{d},\mathbb{C})$. Let $\mathcal{W}$ be a weight function matrix and consider the weighted $L^{\infty}$-type space by
$\mathcal{L}^{\infty}_{\omega^{\ell}}:=\left\{f\in L^{\infty}:\;\;\;\|f\|_{\infty,\omega^{\ell}}:=\operatorname{ess}\sup_{x\in\mathbb{R}^{d}}|f(x)|e^{\omega^{\ell}(x)}<+\infty\right\},$
and define
(3.2) $\mathcal{L}^{\infty}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{L}^{\infty}_{\omega^{\ell}},\hskip 15.0pt\mathcal{L}^{\infty}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{L}^{\infty}_{\omega^{\ell}},$
endowed with their natural topologies. Moreover, for any $1\leq p<\infty$ we assume that $\mathcal{W}$ is continuous and for all (parameters) $\ell>0$ let us set
$\mathcal{L}^{p}_{\omega^{\ell}}:=\left\{f\in L^{p}:\;\;\;\|f\|_{p,\omega^{\ell}}:=\left(\int_{\mathbb{R}^{d}}|f(x)|^{p}e^{\omega^{\ell}(x)}dx\right)^{1/p}<+\infty\right\},$
and put
(3.3) $\mathcal{L}^{p}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{L}^{p}_{\omega^{\ell}},\hskip 15.0pt\mathcal{L}^{p}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{L}^{p}_{\omega^{\ell}}.$
Again, the spaces $\mathcal{L}^{p}_{\{\mathcal{W}\}}$ and $\mathcal{L}^{p}_{(\mathcal{W})}$ are endowed with their natural topologies. Here and throughout the whole paper to lighten notation we use the convention to write $f$ instead of $[f]$; i.e. identify a representative of the equivalence class with the whole class.
Moreover, from now on write $\cdot$] as a joint notation if we mean either $\{\cdot\}$ or $(\cdot)$. A similar convention is also applied to the appearing growth conditions for weight function matrices in the next sections. $\mathcal{L}^{p}_{\{\mathcal{W}\}}$ are called classes of Roumieu-type and $\mathcal{L}^{p}_{(\mathcal{W})}$ of Beurling-type. Note that for the Beurling-type all large parameters $\ell>0$ are becoming crucial, whereas for the Roumieu-type all small parameters $\ell$ are important.
The following obvious and continuous inclusions are valid for any (continuous) weight function matrix:
$\mathcal{L}^{p}_{(\mathcal{W})}\subseteq\mathcal{L}^{p}_{\{\mathcal{W}\}}\subseteq L^{p},\;\;\;1\leq p\leq\infty.$
Note that for any simple weight function matrix the Roumieu- and the Beurling-type coincide. On the other hand, if $\mathcal{W}=\{\omega^{\ell}>0\}$ stabilizes above, i.e. there exists $\ell_{0}>0$ such that $\omega^{\ell}\equiv\omega^{\ell_{0}}$ for all $\ell\geq\ell_{0}$, then $\mathcal{L}^{p}_{(\mathcal{W})}=\mathcal{L}^{p}_{(\omega^{\ell_{0}})}$ when identifying $\omega^{\ell_{0}}$ with the simple matrix $\mathcal{W}=\{\omega^{\ell_{0}}\}$. Similarly, if $\mathcal{W}$ stabilizes below, i.e. there exists $\ell_{0}>0$ such that $\omega^{\ell}\equiv\omega^{\ell_{0}}$ for all $\ell\leq\ell_{0}$, then $\mathcal{L}^{p}_{\{\mathcal{W}\}}=\mathcal{L}^{p}_{\{\omega^{\ell_{0}}\}}$. Therefore, in these cases one can omit all superfluous weight functions $\omega^{\ell}$, $\ell\neq\ell_{0}$, but the Roumieu- and the Beurling-type will not coincide in general.
3.2. On the unboundedness of weight function matrices
Assume that $\omega^{\ell_{0}}=0$ for some $\ell_{0}>0$, then $L^{p}$ coincides as set and l.c.v.s. with $\mathcal{L}^{p}_{\{\mathcal{W}\}}$ and, similarly, when $\omega^{\ell}=0$ for all $\ell>0$ then $L^{p}$ coincides with $\mathcal{L}^{p}_{(\mathcal{W})}$ (and with $\mathcal{L}^{p}_{\{\mathcal{W}\}}$).
More generally, if
(3.4) $\exists\;\ell_{0}>0:\;\;\;\sup_{x\in\mathbb{R}^{d}}\omega^{\ell_{0}}(x)<+\infty,$
then $\mathcal{L}^{p}_{\{\mathcal{W}\}}=L^{p}$ (as sets and l.c.v.s.) and, analogously, if
(3.5) $\forall\;\ell>0:\;\;\;\sup_{x\in\mathbb{R}^{d}}\omega^{\ell}(x)<+\infty,$
then $\mathcal{L}^{p}_{(\mathcal{W})}=L^{p}$ (as sets and l.c.v.s.), $1\leq p\leq\infty$. Now let us prove the converse implication.
Proposition 3.2.
Let $\mathcal{W}=\{\omega^{\ell}:\ell>0\}$ be a weight function matrix.
  • $(i)$
    The inclusion (as sets) $L^{\infty}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{W}\}}$ implies (3.4) whereas $L^{\infty}\subseteq\mathcal{L}^{\infty}_{(\mathcal{W})}$ implies (3.5) for $\mathcal{W}$.
  • $(ii)$
    If $\mathcal{W}$ is continuous, then the inclusion (as sets) $L^{p}\subseteq\mathcal{L}^{p}_{\{\mathcal{W}\}}$ for some/any $1\leq p<\infty$ implies (3.4) and $L^{p}\subseteq\mathcal{L}^{p}_{(\mathcal{W})}$ implies (3.5) for $\mathcal{W}$.
Proof. $(i)$ By assumption the equality $L^{\infty}=\mathcal{L}^{\infty}_{[\mathcal{W}]}$ holds (as sets) and we show that this already implies (3.4) resp. (3.5) for $\mathcal{W}$: Otherwise, take the constant function $f\equiv 1$, and then, if (3.4) resp. (3.5) is violated, we get $\sup_{x\in\mathbb{R}^{d}}|f(x)|e^{\omega^{\ell}(x)}=+\infty$ for all $\ell>0$ resp. for all sufficiently large $\ell>0$. But this implies $f\notin\mathcal{L}^{\infty}_{[\mathcal{W}]}=L^{\infty}$, a contradiction.
$(ii)$ By assumption $L^{p}=\mathcal{L}^{p}_{[\mathcal{W}]}$ holds (as sets) and again we proceed by contradiction.
Assume now that (3.4) is violated for $\mathcal{W}$. Then we can find a sequence $(x_{n})_{n\in\mathbb{N}_{>0}}$ in $\mathbb{R}^{d}$ such that $\omega^{1/n}(x_{n})>n^{2}$ for all $n\in\mathbb{N}_{>0}$. By the continuity of each weight function and the pointwise order of all weight functions in the matrix one can assume that $|x_{n}|\rightarrow+\infty$ and, moreover, the continuity of (each) $\omega^{1/n}$ also implies that
(3.6) $\forall\;n\in\mathbb{N}_{>0}\;\exists\;\delta_{n}>0\;\forall\;x\in I_{n}:=\{y\in\mathbb{R}^{d}:|y-x_{n}|\leq\delta_{n}\}:\;\;\;\omega^{1/n}(x)\geq n^{2}.$
W.l.o.g. we can assume $\delta_{n}\leq 1$ and that the sets (balls) $I_{n}$ are pairwise disjoint. Then set $J_{n}:=\int_{x\in\mathbb{R}^{d}:x\in I_{n}}1dx$ and let $g^{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be defined as follows:
(3.7) $g^{p}(x):=\left(\frac{1}{2^{n}J_{n}}\right)^{1/p},\;\;\;\text{for}\;x\in I_{n},\hskip 15.0ptg^{p}(x):=0,\;\;\;\text{for}\;x\notin I_{n}.$
First, $g^{p}\in L^{p}$ follows because $g^{p}$ is continuous on $\mathbb{R}^{d}$ almost everywhere and
$\int_{\mathbb{R}^{d}}|g^{p}(x)|^{p}dx=\sum_{n=1}^{+\infty}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}|g^{p}(x)|^{p}dx=\sum_{n=1}^{+\infty}\frac{1}{2^{n}J_{n}}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}1dx=\sum_{n=1}^{+\infty}\frac{1}{2^{n}}=1.$
On the other hand let $\ell>0$ be given (small), arbitrary but fixed. There exists $n_{\ell}\in\mathbb{N}_{>0}$ such that $\frac{1}{n}\leq\ell$ for all $n\geq n_{\ell}$ and so, by (3.6) and (3.7) and the pointwise order of the weight functions, for all $n\in\mathbb{N}_{>0}$ with $n\geq n_{\ell}$ and $x\in\mathbb{R}^{d}$ with $x\in I_{n}$ we get $|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}\geq|g^{p}(x)|^{p}e^{\omega^{1/n}(x)}\geq\frac{1}{2^{n}J_{n}}e^{n^{2}}$. Therefore, we can estimate as follows:
$\int_{\mathbb{R}^{d}}|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}dx=\sum_{n=1}^{+\infty}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}dx$
$=\sum_{n=1}^{n_{\ell}-1}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}dx+\sum_{n=n_{\ell}}^{+\infty}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}|g(x)|^{p}e^{\omega^{\ell}(x)}dx$
$\geq\sum_{n=1}^{n_{\ell}-1}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}dx+\sum_{n=n_{\ell}}^{+\infty}\frac{1}{2^{n}J_{n}}e^{n^{2}}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}1dx$
$=\sum_{n=1}^{n_{\ell}-1}\int_{x\in\mathbb{R}^{d}:x\in I_{n}}|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}dx+\sum_{n=n_{\ell}}^{+\infty}\frac{e^{n^{2}}}{2^{n}}=+\infty.$
Since $\ell>0$ was arbitrary, this computation verifies $g^{p}\notin\mathcal{L}^{p}_{\{\mathcal{W}\}}=L^{p}$, a contradiction.
If (3.5) is violated, then there exists some $\ell_{0}>0$ such that $\sup_{x\in\mathbb{R}^{d}}\omega^{\ell}(x)=+\infty$ for all $\ell\geq\ell_{0}$ and so we can find a sequence $(x_{n})_{n\in\mathbb{N}_{>0}}$ in $\mathbb{R}^{d}$ with $|x_{n}|\rightarrow+\infty$ and such that $\omega^{\ell_{0}+1/n}(x_{n})>n^{2}$ for all $n\in\mathbb{N}_{>0}$. Moreover, analogously to (3.6), we get
(3.8) $\forall\;n\in\mathbb{N}_{>0}\;\exists\;\delta_{n}>0\;\forall\;x\in\mathbb{R}^{d},\;x\in I_{n}:=\{y\in\mathbb{R}^{d}:|y-x_{n}|\leq\delta_{n}\}:\;\;\;\omega^{\ell_{0}+1/n}(x)\geq n^{2}.$
Again w.l.o.g. assume $\delta_{n}\leq 1$ and that the sets $I_{n}$ are pairwise disjoint. Let $g^{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be defined as in (3.7) and so $g^{p}\in L^{p}$ follows as before. Now let $\ell>\ell_{0}$ be given (large), arbitrary but fixed. There exists $n_{\ell}\in\mathbb{N}_{>0}$ such that $\ell_{0}+\frac{1}{n}\leq\ell$ for all $n\geq n_{\ell}$ and so, by (3.8), for all $n\in\mathbb{N}_{>0}$ with $n\geq n_{\ell}$ and $x\in\mathbb{R}^{d}$ with $x\in I_{n}$ we get $|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}\geq|g^{p}(x)|^{p}e^{\omega^{\ell_{0}+1/n}(x)}\geq\frac{1}{2^{n}J_{n}}e^{n^{2}}$. Using this and the same estimate of the crucial integral as before one infers
$\int_{\mathbb{R}^{d}}|g^{p}(x)|^{p}e^{\omega^{\ell}(x)}dx=+\infty,$
and so, since $\ell>\ell_{0}$ was arbitrary, $g^{p}\notin\mathcal{L}^{p}_{(\mathcal{W})}$ which yields a contradiction. ∎
Proposition 3.2 and the previous comments immediately imply the following:
Theorem 3.3.
Let $\mathcal{W}=\{\omega^{\ell}:\ell>0\}$ be a weight function matrix.
  • $(i)$
    (3.4) holds if and only if $L^{\infty}=\mathcal{L}^{\infty}_{\{\mathcal{W}\}}$ as sets and/or l.c.v.s. and (3.5) if and only if $L^{\infty}=\mathcal{L}^{\infty}_{(\mathcal{W})}$ as sets and/or l.c.v.s.
  • $(ii)$
    If $\mathcal{W}$ is continuous, then (3.4) holds if and only if $L^{p}=\mathcal{L}^{p}_{\{\mathcal{W}\}}$ and (3.5) if and only if $L^{p}=\mathcal{L}^{p}_{(\mathcal{W})}$ and the equality holds for some/any $1\leq p<\infty$ as sets and/or l.c.v.s.
Therefore, in order to treat proper (new) weighted subclasses of $L^{p}$ it is natural to assume always that $\mathcal{W}$ is unbounded; i.e. that (3.1) holds. Note that in the Beurling-type case, if (3.5) formally fails and (3.1) is (only) valid for all $\ell\geq\ell_{0}$ (or $\ell>\ell_{0}$) large, then by definition of the spaces we can skip all $0<\ell<\ell_{0}$ (or $0<\ell\leq\ell_{0}$) without changing the weighted class since we are interested in all large $\ell$. Hence (3.1) is natural for both types and analogous comments explain the notions for $\mathcal{W}$ being radial, non-decreasing or continuous.
However, note that in the forthcoming proofs it is not required necessarily that the weight function matrices are unbounded. We finish this section by commenting on the (formal) opposite extreme.
Remark 3.4.
Let $\mathcal{W}:=\{\omega^{\ell}:\ell>0\}$ be given and assume that there exists $\ell_{0}>0$ such that $\omega^{\ell}\equiv+\infty$ for all $\ell\geq\ell_{0}$ resp. such that $\omega^{\ell}\equiv+\infty$ for all $\ell>0$ (this notion means $\omega^{\ell}(x)=+\infty$ for all $x\in\mathbb{R}^{d}$). In this case, formally, it makes sense to set $\mathcal{L}^{p}_{(\mathcal{W})}=\{0\}$ resp. $\mathcal{L}^{p}_{\{\mathcal{W}\}}=\{0\}$, $1\leq p\leq\infty$, and this convention does not change when (only) assuming that for any $\ell\geq\ell_{0}$ (resp. any $\ell>0$) one can find some $x_{\ell}\in\mathbb{R}^{d}$ such that $\omega^{\ell}(x)=+\infty$ for all $x\in\mathbb{R}^{d}$ such that $|x|\geq|x_{\ell}|$.
Conversely, we use the convention to identify $\{0\}$ with the weighted $L^{p}$-type spaces when being formally defined via the previously described “non-standard weight matrices”.
3.3. On the non-triviality of weighted $L^{p}$-type spaces
We verify that the defined weighted spaces are non-trivial under mild and natural assumptions on the (radial) weight function matrix $\mathcal{W}$.
Lemma 3.5.
Let $\mathcal{W}$ be non-decreasing and radial.
  • $(i)$
    It holds that $\{0\}\subsetneq\mathcal{L}^{\infty}_{(\mathcal{W})}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{W}\}}$.
  • $(ii)$
    If $\mathcal{W}$ is in addition continuous, then $\{0\}\subsetneq\mathcal{L}^{p}_{(\mathcal{W})}\subseteq\mathcal{L}^{p}_{\{\mathcal{W}\}}$ for all $1\leq p<\infty$.
Proof. $(i)$ If (3.5) is valid, then $\mathcal{L}^{\infty}_{(\mathcal{W})}=L^{\infty}\neq\{0\}$ by Theorem 3.3.
If (3.5) is violated and since $\mathcal{W}$ is non-decreasing we can find $\ell_{0}>0$ such that $\lim_{|x|\rightarrow\infty}\omega^{\ell}(x)=+\infty$ for all $\ell\geq\ell_{0}$.
We introduce the following functions:
$\sigma^{\infty}(x):=\omega^{n+1}(x),\;\;\;\forall\;x\in\mathbb{R}^{d},\;|x|\in[n,n+1),\;\;\;n\in\mathbb{N},$
and
(3.9) $\psi^{\infty}_{\mathcal{W}}(x):=e^{-\sigma^{\infty}(x)^{2}}.$
So for any $\ell>0$ and all $x\in\mathbb{R}^{d}$ such that $|x|\in[n,n+1)$ one has $|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}=e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{2}}$. Let $\ell\geq\ell_{0}$ be fixed and hence for all $|x|\in[n,n+1)$ with $n\geq\ell$ we have $|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}\leq e^{\omega^{\ell}(x)(1-\omega^{\ell}(x))}$ and this last expression tends to 0 as $|x|\rightarrow+\infty$. Thus,
$\forall\;\ell\geq\ell_{0}\;\exists\;C_{\ell}\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;|\psi^{\infty}_{\mathcal{W}}(x)|e^{\omega^{\ell}(x)}\leq C_{\ell},$
and so $\psi^{\infty}_{\mathcal{W}}\in\mathcal{L}^{\infty}_{(\mathcal{W})}$.
$(ii)$ If (3.5) holds, then $\mathcal{L}^{p}_{(\mathcal{W})}=L^{p}\neq\{0\}$ by Theorem 3.3.
If (3.5) fails, then since $\mathcal{W}$ is non-decreasing there exists $\ell_{0}>0$ such that $\lim_{|x|\rightarrow\infty}\omega^{\ell}(x)=+\infty$ for all $\ell\geq\ell_{0}$. In particular, there exists $x_{0}\in\mathbb{R}^{d}$, $|x_{0}|\geq 1$, such that $\omega^{\ell}(x)>1$ for all $\ell\geq\ell_{0}$ and all $x\in\mathbb{R}^{d}$ with $|x|\geq|x_{0}|$. W.l.o.g. assume that $|x_{0}|=n_{0}=\ell_{0}\in\mathbb{N}_{>0}$. Now, for any $n\in\mathbb{N}_{>0}$ with $n\geq n_{0}$ we can find real numbers $a_{n}\geq 1$ such that
(3.10) $(\omega^{n_{0}}(x))^{a_{n}}\geq 2(d+1)\log(1+|x|),\;\;\;\forall\;x\in\mathbb{R}^{d},\;|x|\in[n,n+1),\;n\geq n_{0}.$
Indeed, for this it suffices to take $a_{n}:=\max\{1,\frac{\log(2(d+1)\log(2+n))}{\log(\omega^{n_{0}}(n))}\}$ and (3.10) holds as well for all $\omega^{n}$, $n\geq n_{0}$. Moreover, for all $n\geq n_{0}$ choose positive reals $b_{n}\geq 1$ such that
(3.11) $(\omega^{n}(x))^{b_{n}}\geq 2\omega^{n}(x),\;\;\;\forall\;x\in\mathbb{R}^{d},\;|x|\in[n,n+1),\;n\geq n_{0}.$
When enlarging $a_{n}$, if necessary, it suffices to take
$b_{n}=\frac{a_{n}\log(\omega^{n}(n+1))+\log(2)}{\log(\omega^{n}(n))}(\geq\max\{1,\frac{\log(\omega^{n}(n+1))+\log(2)}{\log(\omega^{n}(n))}\}).$
Finally, in order to complete the definition one can set $a_{n}:=1$ and $b_{n}:=1$ for $0\leq n<n_{0}$.
Now introduce the auxiliary weight function
$\sigma^{p}(x):=(\omega^{n+1}(x))^{a_{n}b_{n}},\;\;\;\forall\;x\in\mathbb{R}^{d},\;|x|\in[n,n+1),\;n\in\mathbb{N},$
and set
(3.12) $\psi^{p}_{\mathcal{W}}(x):=e^{-p^{-1}\sigma^{p}(x)}.$
Then, for any $\ell\geq\ell_{0}$ (fixed) we estimate as follows:
$\int_{\mathbb{R}^{d}}|\psi^{p}_{\mathcal{W}}(x)|^{p}e^{\omega^{\ell}(x)}dx=\sum_{n=0}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx$
$=\sum_{n=0}^{\ell-1}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n=\ell}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx$
$\leq\sum_{n=0}^{\ell-1}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n=\ell}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{n}(x)-(\omega^{n}(x))^{a_{n}b_{n}}}dx$
$\leq\sum_{n=0}^{\ell-1}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}e^{\omega^{\ell}(x)-(\omega^{n+1}(x))^{a_{n}b_{n}}}dx+\sum_{n=\ell}^{+\infty}\int_{x\in\mathbb{R}^{d},\;|x|\in[n,n+1)}\frac{1}{(1+|x|^{2})^{d+1}}dx<+\infty.$
The last inequality is valid since $\frac{1}{(1+|x|)^{2(d+1)}}\leq\frac{1}{(1+|x|^{2})^{(d+1)}}$ for all $x\in\mathbb{R}^{d}$. For the second last one note that $e^{\omega^{n}(x)-(\omega^{n}(x))^{a_{n}b_{n}}}\leq\frac{1}{(1+|x|)^{2(d+1)}}$ holds on the intervals under consideration because for $|x|\in[n,n+1)$, $n\geq\ell_{0}=n_{0}$, via the choices of $a_{n}$ and $b_{n}$ in (3.10) and (3.11):
$(\omega^{n}(x))^{a_{n}b_{n}}-\omega^{n}(x)\geq(\omega^{n}(x))^{a_{n}b_{n}}-(\omega^{n}(x))^{a_{n}}\geq(\omega^{n}(x))^{a_{n}}\geq(\omega^{\ell_{0}}(x))^{a_{n}}\geq 2(d+1)\log(1+|x|).$
Consequently, since $\ell\geq\ell_{0}$ was arbitrary it follows that $\psi^{p}_{\mathcal{W}}\in\mathcal{L}^{p}_{(\mathcal{W})}$. ∎
3.4. Stability under translation
For concrete applications it is important to know whether weighted spaces satisfy certain stability properties. Concerning the classes introduced above we investigate invariance w.r.t. translations. This property is somehow natural and crucially required in the proofs of the main characterizing results in the next section.
Notation: For given $f:\mathbb{R}^{d}\rightarrow\mathbb{C}$ and $x_{0}\in\mathbb{R}^{d}$ set $f_{x_{0}}:=f(x-x_{0})$ and consider the translation operator $T_{x_{0}}:f\mapsto f_{x_{0}}$.
The main statement in this context is the following result:
Proposition 3.6.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices.
  • $(i)$
    If the matrices are related by
    (3.13) $\forall\;\ell>0\;\exists\;n>0\;\exists\;L\geq 1\;\forall\;x,y\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x+y)\leq\sigma^{n}(x)+\sigma^{n}(y)+L,$
    then
    $\forall\;\ell>0\;\exists\;n>0\;\exists\;L\geq 1\;\forall\;x_{0}\in\mathbb{R}^{d}\;\forall\;f\in\mathcal{L}^{\infty}_{\sigma^{n}}:\;\;\;\|T_{x_{0}}(f)\|_{\infty,\tau^{\ell}}\leq e^{L}e^{\sigma^{n}(x_{0})}\|f\|_{\infty,\sigma^{n}}.$
    If both matrices are in addition continuous, then for any $1\leq p<\infty$ we get
    $\forall\;\ell>0\;\exists\;n>0\;\exists\;L\geq 1\;\forall\;x_{0}\in\mathbb{R}^{d}\;\forall\;f\in\mathcal{L}^{p}_{\sigma^{n}}:\;\;\;\|T_{x_{0}}(f)\|^{p}_{p,\tau^{\ell}}\leq e^{L}e^{\sigma^{n}(x_{0})}\|f\|^{p}_{p,\sigma^{n}}.$
  • $(ii)$
    If the matrices are related by
    (3.14) $\forall\;n>0\;\exists\;\ell>0\;\exists\;L\geq 1\;\forall\;x,y\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x+y)\leq\sigma^{n}(x)+\sigma^{n}(y)+L,$
    then
    $\forall\;n>0\;\exists\;\ell>0\;\exists\;L\geq 1\;\forall\;x_{0}\in\mathbb{R}^{d}\;\forall\;f\in\mathcal{L}^{\infty}_{\sigma^{n}}:\;\;\;\|T_{x_{0}}(f)\|_{\infty,\tau^{\ell}}\leq e^{L}e^{\sigma^{n}(x_{0})}\|f\|_{\infty,\sigma^{n}}.$
If both matrices are in addition continuous, then for any $1\leq p<\infty$ we get
$\forall\;n>0\;\exists\;\ell>0\;\exists\;L\geq 1\;\forall\;x_{0}\in\mathbb{R}^{d}\;\forall\;f\in\mathcal{L}^{p}_{\sigma^{n}}:\;\;\;\|T_{x_{0}}(f)\|^{p}_{p,\tau^{\ell}}\leq e^{L}e^{\sigma^{n}(x_{0})}\|f\|^{p}_{p,\sigma^{n}}.$
Proposition 3.6 implies the continuity of (each) operator $T_{x_{0}}$ between weighted $L^{p}$-type classes and note:
    The choice for the new index $n$ is only depending on given $\ell$ via (3.13) resp. the index $\ell$ is depending on given $n$ via (3.14) and hence not on the translation parameter $x_{0}$. Moreover, it is also not depending on the choice of the representative of the equivalence class.
    The statement becomes unclear if $\mathbb{R}^{d}$ is replaced by some open and bounded $U\subseteq\mathbb{R}^{d}$.
    The special case $\mathcal{S}=\mathcal{T}$, i.e. (3.13) for $\mathcal{S}$, implies that the classes $\mathcal{L}^{p}_{(\mathcal{S})}$, $1\leq p\leq\infty$, are translation invariant and each $T_{x_{0}}$ is acting continuously on the corresponding weighted space. And (3.14) for $\mathcal{S}$ implies that $\mathcal{L}^{p}_{\{\mathcal{S}\}}$, $1\leq p\leq\infty$, is translation invariant and each $T_{x_{0}}$ is acting continuously.
Proof. We only prove $(i)$ in detail since $(ii)$ is analogous. When taking into account (3.13) we get for all $x_{0}\in\mathbb{R}^{d}$:
$\|T_{x_{0}}(f)\|_{\infty,\tau^{\ell}}$ $=\|f_{x_{0}}\|_{\infty,\tau^{\ell}}=\operatorname{ess}\sup_{x\in\mathbb{R}^{d}}|f_{x_{0}}(x)|e^{\tau^{\ell}(x)}=\operatorname{ess}\sup_{x\in\mathbb{R}^{d}}|f(x-x_{0})|e^{\tau^{\ell}(x)}$
$=\operatorname{ess}\sup_{y\in\mathbb{R}^{d}}|f(y)|e^{\tau^{\ell}(y+x_{0})}\leq e^{L}e^{\sigma^{n}(x_{0})}\operatorname{ess}\sup_{y\in\mathbb{R}^{d}}|f(y)|e^{\sigma^{n}(y)}=e^{L}e^{\sigma^{n}(x_{0})}\|f\|_{\infty,\sigma^{n}}.$
Similarly, for all $1\leq p<\infty$ and $x_{0}\in\mathbb{R}^{d}$:
$\|T_{x_{0}}(f)\|^{p}_{p,\tau^{\ell}}$ $=\|f_{x_{0}}\|^{p}_{p,\tau^{\ell}}=\int_{\mathbb{R}^{d}}|f_{x_{0}}(x)|^{p}e^{\tau^{\ell}(x)}dx=\int_{\mathbb{R}^{d}}|f(x-x_{0})|^{p}e^{\tau^{\ell}(x)}dx$
$=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\tau^{\ell}(y+x_{0})}dy\leq e^{L}e^{\sigma^{n}(x_{0})}\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\sigma^{n}(y)}dy=e^{L}e^{\sigma^{n}(x_{0})}\|f\|^{p}_{p,\sigma^{n}}.$
We continue with an auxiliary technical result for radial weight matrices which should be compared with $(f)$ in Section 2.4.
Lemma 3.7.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be non-decreasing and radial weight function matrices. Consider the following (mixed) assertions:
  • $(i)$
    The matrices are related by
    (3.15) $\forall\;\ell>0\;\exists\;n>0\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\tau^{\ell}(2t)\leq\sigma^{n}(t)+L.$
  • $(ii)$
    The matrices are related by (3.13).
  • $(iii)$
    The matrices are related by
    (3.16) $\forall\;\ell>0\;\exists\;n>0\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\tau^{\ell}(2t)\leq 2\sigma^{n}(t)+L.$
Then the implications $(i)\Rightarrow(ii)\Rightarrow(iii)$ are valid. The proof shows that we can choose for given $\ell$ in each assertion the same parameter $n$ and the assumption “non-decreasing” is only used to show the first implication. The analogous statement holds when $\forall\;\ell>0\;\exists\;n>0$” is replaced by $\forall\;n>0\;\exists\;\ell>0$” and involving (3.14); for future reference we only mention explicitly
(3.17) $\forall\;n>0\;\exists\;\ell>0\;\exists\;L\geq 1\;\forall\;t\geq 0:\;\;\;\tau^{\ell}(2t)\leq\sigma^{n}(t)+L.$
Proof. Since $\tau^{\ell}$ and $\sigma^{n}$ are non-decreasing, by (3.15) one has $\tau^{\ell}(s+t)\leq\tau^{\ell}(2\max\{s,t\})\leq\sigma^{n}(\max\{s,t\})+L\leq\sigma^{n}(s)+\sigma^{n}(t)+L$ for all $s,t\geq 0$. On the other hand, in (3.13) we set $s=t$ and immediately get (3.16). The second part follows analogously. ∎
This technical lemma motivates the next comment:
Remark 3.8.
In view of Lemma 3.7 the conclusions in Proposition 3.6 hold when assuming for non-decreasing, radial (and continuous) matrices $\mathcal{S}$, $\mathcal{T}$, properties (3.15) resp. (3.17) which are “more natural” and easier to check.
Instead of (3.15) one could expect the weaker mixed $(\omega_{1})$-like condition
(3.18) $\forall\;\ell>0\;\exists\;n>0:\;\;\;\tau^{\ell}(2t)=O(\sigma^{n}(t))\;\;\;t\rightarrow+\infty,$
and instead of (3.17) the condition
(3.19) $\forall\;n>0\;\exists\;\ell>0:\;\;\;\tau^{\ell}(2t)=O(\sigma^{n}(t))\;\;\;t\rightarrow+\infty.$
When both matrices are non-decreasing, then (3.18) is equivalent to $\tau^{\ell}(s+t)\leq L(\sigma^{n}(s)+\sigma^{n}(t))+L$ for some $L\geq 1$ and all $s,t\geq 0$ and similarly for (3.19); the relation between the indices $\ell$ and $n$ is preserved in this equivalence. This follows by generalizing directly the arguments in $(f)$ in Section 2.4 to the mixed setting. However, when involving (3.18) and (3.19) the proof of Proposition 3.6 yields terms $e^{L\sigma^{n}(y)}$. By the appearance of the additional multiplicative constant $L$ in the exponent the conclusion in Proposition 3.6 becomes unclear in general. Similar comments apply to the growth relations introduced in the next section; we refer to Section 5.1 for more explanations.
4. Characterization of inclusion relations for weighted $L^{p}$-type spaces
The aim is to prove the characterization of inclusion relations between weighted $L^{p}$-type spaces defined by weight function matrices when assuming mild standard growth requirements. In order to proceed, first let us introduce the crucial growth relations. Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$, $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be given and consider
(4.1) $\forall\;\ell>0\;\exists\;n>0\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{n}(x)+C,$
and
(4.2) $\forall\;n>0\;\exists\;\ell>0\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{n}(x)+C.$
We write $\mathcal{S}(\preceq)\mathcal{T}$ if (4.1) is valid and $\mathcal{S}\{\preceq\}\mathcal{T}$ if (4.2) holds and using this notation the goal is to verify the following main characterizing statements:
Theorem 4.1.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy (3.13). Then the following are equivalent:
  • $(i)$
    The matrices are related by $\mathcal{S}(\preceq)\mathcal{T}$.
  • $(ii)$
    The inclusion $\mathcal{L}^{\infty}_{(\mathcal{S})}\subseteq\mathcal{L}^{\infty}_{(\mathcal{T})}$ holds as sets.
  • $(iii)$
    The continuous inclusion $\mathcal{L}^{\infty}_{(\mathcal{S})}\subseteq\mathcal{L}^{\infty}_{(\mathcal{T})}$ holds.
If both matrices are also continuous, then analogous equivalences are valid when the symbol/functor $\mathcal{L}^{\infty}$ is replaced by $\mathcal{L}^{p}$ for some/any $1\leq p<\infty$.
Theorem 4.2.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy (3.14). Then the following are equivalent:
  • $(i)$
    The matrices are related by $\mathcal{S}\{\preceq\}\mathcal{T}$.
  • $(ii)$
    The inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ holds as sets.
  • $(iii)$
    The continuous inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ holds.
If both matrices are also continuous, then analogous equivalences are valid when the symbol/functor $\mathcal{L}^{\infty}$ is replaced by $\mathcal{L}^{p}$ for some/any $1\leq p<\infty$.
First, the next statement is immediate by the defining system of seminorms and the introduced growth relations before and it verifies $(i)\Rightarrow(ii),(iii)$ in Theorems 4.1 and 4.2.
Proposition 4.3.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices.
  • $(i)$
    If $\mathcal{S}(\preceq)\mathcal{T}$, then
    $\forall\;\ell>0\;\exists\;n>0\;\exists\;C\geq 1\;\forall\;f\in\mathcal{L}^{\infty}_{\sigma^{n}}:\;\;\;\|f\|_{\infty,\tau^{\ell}}\leq e^{C}\|f\|_{\infty,\sigma^{n}},$
    and if $\mathcal{S}\{\preceq\}\mathcal{T}$, then
    $\forall\;n>0\;\exists\;\ell>0\;\exists\;C\geq 1\;\forall\;f\in\mathcal{L}^{\infty}_{\sigma^{n}}:\;\;\;\|f\|_{\infty,\tau^{\ell}}\leq e^{C}\|f\|_{\infty,\sigma^{n}}.$
    Thus $\mathcal{S}[\preceq]\mathcal{T}$ implies $\mathcal{L}^{\infty}_{[\mathcal{S}]}\subseteq\mathcal{L}^{\infty}_{[\mathcal{T}]}$ with continuous inclusion.
  • $(ii)$
    If both $\mathcal{S}$ and $\mathcal{T}$ are in addition continuous, then analogously $\mathcal{S}[\preceq]\mathcal{T}$ implies $\mathcal{L}^{p}_{[\mathcal{S}]}\subseteq\mathcal{L}^{p}_{[\mathcal{T}]}$ with continuous inclusion for some/any $1\leq p<\infty$. (The only difference compared with $(i)$ is that in the estimates $e^{C}$ has to be replaced by $e^{C/p}$.)
The aim now is to verify the (expected) characterization of inclusion relations in terms of $(\preceq)$ and $\{\preceq\}$ between weight function matrices. For this we have to impose mild extra conditions on the weights; in Section 4.1 we present for both the Beurling- and the Roumieu-type an abstract functional analytic argument (see Propositions 4.5 and 4.6). This approach, in particular the Beurling-type result, has been inspired by [3, Thm. 1.3.18]. Afterwards, in Section 4.2, we give for the Roumieu-type a different constructive approach involving special (optimal) functions.
In the proofs of Propositions 4.5 and 4.6 the non-triviality of $\mathcal{L}^{p}_{[\mathcal{S}]}$ is required and in the isotropic setting in view of Lemma 3.5 it suffices to assume in addition that $\mathcal{S}$ is non-decreasing. However, by taking into account the convention from Remark 3.4 we comment now on the fact that, if these spaces are trivial, then the implications follow automatically.
Remark 4.4.
When $\{0\}=\mathcal{L}^{p}_{[\mathcal{S}]}\subseteq\mathcal{L}^{p}_{[\mathcal{T}]}$, $1\leq p\leq\infty$, then the inclusion is trivial and the convention from Remark 3.4 gives that $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ corresponds to the weight function matrix satisfying $\sigma^{\ell}\equiv+\infty$ for all $\ell>0$ sufficiently large in the Beurling-case resp. $\sigma^{\ell}\equiv+\infty$ for any $\ell>0$ in the Roumieu-case. And therefore, by recalling (4.1) resp. (4.2), the relation $\mathcal{S}(\preceq)\mathcal{T}$ resp. $\mathcal{S}\{\preceq\}\mathcal{T}$ is formally trivial.
Finally, by using the convention $\infty$$\infty$, the relations formally even follow when in addition $\mathcal{L}^{p}_{[\mathcal{T}]}=\{0\}$.
4.1. Proof of the main characterizing results
We start with the Beurling-type.
Proposition 4.5.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices and assume that both matrices satisfy (3.13).
  • $(i)$
    If both matrices are also continuous, then the (continuous) inclusion $\mathcal{L}^{p}_{(\mathcal{S})}\subseteq\mathcal{L}^{p}_{(\mathcal{T})}$ (for some/any $1\leq p<\infty$) implies $\mathcal{S}(\preceq)\mathcal{T}$.
  • $(ii)$
    The (continuous) inclusion $\mathcal{L}^{\infty}_{(\mathcal{S})}\subseteq\mathcal{L}^{\infty}_{(\mathcal{T})}$ implies $\mathcal{S}(\preceq)\mathcal{T}$.
Proof. $(i)$ First, by (3.13) we get:
$\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;L\geq 1\;\forall\;y,x_{0}\in\mathbb{R}^{d}:$
$\tau^{\ell}(x_{0})=\tau^{\ell}(x_{0}+y-y)\leq\tau^{\ell^{\prime}}(x_{0}+y)+\tau^{\ell^{\prime}}(-y)+L.$
Thus, for any $f\in\mathcal{L}^{p}_{(\mathcal{S})}\subseteq\mathcal{L}^{p}_{(\mathcal{T})}\subseteq L^{p}$ and $x_{0}\in\mathbb{R}^{d}$ it holds that
$\|f_{x_{0}}\|^{p}_{p,\tau^{\ell^{\prime}}}=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\tau^{\ell^{\prime}}(y+x_{0})}dy\geq e^{-L}e^{\tau^{\ell}(x_{0})}\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{-\tau^{\ell^{\prime}}(-y)}dy,$
and clearly $A_{1}:=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{-\tau^{\ell^{\prime}}(-y)}dy\leq\int_{\mathbb{R}^{d}}|f(y)|^{p}dy<+\infty$ by assumption.
Second, since both $\mathcal{L}^{p}_{(\mathcal{S})}$ and $\mathcal{L}^{p}_{(\mathcal{T})}$ are Fréchet spaces (w.r.t. their natural topologies) we have that the inclusion $\mathcal{L}^{p}_{(\mathcal{S})}\subseteq\mathcal{L}^{p}_{(\mathcal{T})}$ is continuous (in any case); see e.g. [11, Prop. 4.5 & Rem. 4.6]. Consequently,
(4.3) $\forall\;\ell^{\prime}>0\;\exists\;\ell^{\prime\prime}>0\;\exists\;A\geq 1\;\forall\;f\in\mathcal{L}^{p}_{(\mathcal{S})}:\;\;\;\|f\|_{p,\tau^{\ell^{\prime}}}\leq A\|f\|_{p,\sigma^{\ell^{\prime\prime}}}.$
Now, in view of Remark 4.4 we can assume that there exists $0\neq f\in\mathcal{L}^{p}_{(\mathcal{S})}$ and $(i)$ in Proposition 3.6 yields $f_{x_{0}}\in\mathcal{L}^{p}_{(\mathcal{S})}$ for all $x_{0}\in\mathbb{R}^{d}$. The idea is to apply (4.3) to the family $\{f_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$.
Let $\ell>0$ be given, arbitrary but from now on fixed. Let $\ell^{\prime}>0$ be the index appearing in the estimate in the first step and being related to $\ell$ via (3.13). Then apply (4.3) to $\ell^{\prime}$ and let $\ell^{\prime\prime}$ be the index subject to $\ell^{\prime}$ in this property.
Finally, when applying (3.13) to $\mathcal{S}$ and to the index $\ell^{\prime\prime}$ then there exist some $\ell^{\prime\prime\prime}>0$ and $L_{1}\geq 1$ such that for all $x_{0}\in\mathbb{R}^{d}$:
$\|f_{x_{0}}\|^{p}_{p,\sigma^{\ell^{\prime\prime}}}=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\sigma^{\ell^{\prime\prime}}(y+x_{0})}dy\leq e^{L_{1}}e^{\sigma^{\ell^{\prime\prime\prime}}(x_{0})}\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\sigma^{\ell^{\prime\prime\prime}}(y)}dy.$
Note that $A_{2}:=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\sigma^{\ell^{\prime\prime\prime}}(y)}dy<+\infty$ since $f\in\mathcal{L}^{p}_{(\mathcal{S})}$.
Combining everything, so far we have shown
$\forall\;\ell>0\;\exists\;\ell^{\prime\prime\prime}>0\;\exists\;A,L,L_{1}\geq 1\;\exists\;A_{1},A_{2}>0\;\forall\;x_{0}\in\mathbb{R}^{d}:$
$\tau^{\ell}(x_{0})\leq L-\log(A_{1})+p\log(A)+L_{1}+\log(A_{2})+\sigma^{\ell^{\prime\prime\prime}}(x_{0});$
so (4.1) is verified between the indices $\ell$ and $\ell^{\prime\prime\prime}$ and with $C:=L-\log(A_{1})+p\log(A)+L_{1}+\log(A_{2})$; i.e. relation $\mathcal{S}(\preceq)\mathcal{T}$.
$(ii)$ Similarly, since $\mathcal{L}^{\infty}_{(\sigma)}$ and $\mathcal{L}^{\infty}_{(\tau)}$ are Fréchet spaces (w.r.t. their natural topology), by [11, Prop. 4.5 & Rem. 4.6] we have again that the inclusion $\mathcal{L}^{\infty}_{(\mathcal{S})}\subseteq\mathcal{L}^{\infty}_{(\mathcal{T})}$ is continuous in any case. Thus
(4.4) $\forall\;\ell^{\prime}>0\;\exists\;\ell^{\prime\prime}>0\;\exists\;A\geq 1\;\forall\;f\in\mathcal{L}^{\infty}_{(\mathcal{S})}:\;\;\;\|f\|_{\infty,\tau^{\ell^{\prime}}}\leq A\|f\|_{\infty,\sigma^{\ell^{\prime\prime}}},$
and we apply this crucial estimate again to the family $\{f_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$ corresponding to a fixed $0\neq f\in\mathcal{L}^{\infty}_{(\sigma)}$; recall Remark 4.4. The rest follows analogously as in $(i)$ by involving (3.13) and $(i)$ in Proposition 3.6 for both matrices and note that $A_{1}:=\operatorname{ess}\sup_{y\in\mathbb{R}^{d}}|f(y)|e^{-\tau^{\ell}(-y)}\leq\operatorname{ess}\sup_{y\in\mathbb{R}^{d}}|f(y)|<+\infty$ for all $\ell>0$ because $f\in\mathcal{L}^{\infty}_{(\mathcal{T})}$ and $A_{2}:=\operatorname{ess}\sup_{x\in\mathbb{R}^{d}}|f(y)|e^{\sigma^{\ell^{\prime\prime\prime}}(y)}<+\infty$ for any $\ell^{\prime\prime\prime}>0$ because even $f\in\mathcal{L}^{\infty}_{(\mathcal{S})}$ (by assumption). Therefore, $\mathcal{S}(\preceq)\mathcal{T}$ follows as before. ∎
The next result deals with the analogous result for the Roumieu case.
Proposition 4.6.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be given weight function matrices and assume that both matrices satisfy (3.14).
  • $(i)$
    If both matrices are also continuous, then the (continuous) inclusion $\mathcal{L}^{p}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{p}_{\{\mathcal{T}\}}$ for some/any $1\leq p<\infty$ implies $\mathcal{S}\{\preceq\}\mathcal{T}$.
  • $(ii)$
    The (continuous) inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ implies $\mathcal{S}\{\preceq\}\mathcal{T}$.
Proof. $(i)$ We proceed analogously as in Proposition 4.5: By the (continuous) inclusion $\mathcal{L}^{p}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{p}_{\{\mathcal{T}\}}$ and since we are dealing with (LB)-spaces, by de Wilde’s closed graph theorem and Grothendieck’s factorization theorem we obtain
(4.5) $\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;A\geq 1\;\forall\;f\in\mathcal{L}^{p}_{\sigma^{\ell}}:\;\;\;\|f\|_{p,\tau^{\ell^{\prime}}}\leq A\|f\|_{p,\sigma^{\ell}}.$
Fix an index $\ell_{0}>0$ and $0\neq f\in\mathcal{L}^{p}_{\sigma^{\ell_{0}}}$ (recall Remark 4.4). Moreover, assumption (3.14) for $\mathcal{S}$ and $(ii)$ in Proposition 3.6 yield the existence of some $\ell>0$ such that $f_{x_{0}}\in\mathcal{L}^{p}_{\sigma^{\ell}}$ for any $x_{0}\in\mathbb{R}^{d}$. (The indices $\ell_{0}$ and $\ell$ are related by (3.14).) Indeed, the proof of $(ii)$ in Proposition 3.6 gives
$\|f_{x_{0}}\|^{p}_{p,\sigma^{\ell}}\leq e^{L}e^{\sigma^{\ell_{0}}(x_{0})}\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\sigma^{\ell_{0}}(y)}dy.$
Note that $A_{1}:=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\sigma^{\ell_{0}}(y)}dy<+\infty$ since $f\in\mathcal{L}^{p}_{\sigma^{\ell_{0}}}$. We apply (4.5) to the family $\{f_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$ and to this index $\ell$. Moreover, when using (3.14) for $\mathcal{T}$, then analogously as in the proof above for some index $\ell^{\prime}$ (being related to $\ell$ via (4.5)) we can find some index $\ell^{\prime\prime}>0$ and $L_{1}\geq 1$ and get for all $x_{0}\in\mathbb{R}^{d}$:
$\|f_{x_{0}}\|^{p}_{p,\tau^{\ell^{\prime}}}=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{\tau^{\ell^{\prime}}(y+x_{0})}dy\geq e^{-L_{1}}e^{\tau^{\ell^{\prime\prime}}(x_{0})}\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{-\tau^{\ell^{\prime}}(-y)}dy.$
Clearly $A_{2}:=\int_{\mathbb{R}^{d}}|f(y)|^{p}e^{-\tau^{\ell^{\prime}}(-y)}dy\leq\int_{\mathbb{R}^{d}}|f(y)|^{p}dy<+\infty$ since $f\in\mathcal{L}^{p}_{\{\mathcal{T}\}}$ (by assumption). Combining everything, so far we have shown
$\forall\;\ell_{0}>0\;\exists\;\ell^{\prime\prime}>0\,\exists\;A,L,L_{1}\geq 1\;\exists\;A_{1},A_{2}>0\;\forall\;x_{0}\in\mathbb{R}^{d}:$
$\tau^{\ell^{\prime\prime}}(x_{0})\leq L_{1}-\log(A_{2})+p\log(A)+L+\log(A_{1})+\sigma^{\ell_{0}}(x_{0}),$
which yields (4.2) and so $\mathcal{S}\{\preceq\}\mathcal{T}$.
$(ii)$ Similarly, by the continuous inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ and since we are dealing with (LB)-spaces we obtain
(4.6) $\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;A\geq 1\;\forall\;f\in\mathcal{L}^{\infty}_{\sigma^{\ell}}:\;\;\;\|f\|_{\infty,\tau^{\ell^{\prime}}}\leq A\|f\|_{\infty,\sigma^{\ell}}.$
Again we fix an index $\ell_{0}>0$ and $0\neq f\in\mathcal{L}^{\infty}_{\sigma^{\ell_{0}}}$ (recall Remark 4.4) and $(ii)$ in Proposition 3.6 yields the existence of some $\ell>0$ such that $f_{x_{0}}\in\mathcal{L}^{p}_{\sigma^{\ell}}$ for any $x_{0}\in\mathbb{R}^{d}$. (Again the indices $\ell_{0}$ and $\ell$ are related by (3.14).)
We apply (4.6) to the family $\{f_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$ and to this index $\ell$. Then follow the estimates in $(i)$ using the same notation (for the indices) and by involving $(ii)$ in Proposition 3.6 and (3.14); see also $(ii)$ in Proposition 4.5. Therefore $\mathcal{S}\{\preceq\}\mathcal{T}$ holds and note that here $A_{1}:=\operatorname{ess}\sup_{x\in\mathbb{R}^{d}}|f(y)|e^{\sigma^{\ell_{0}}(y)}<+\infty$ because $f\in\mathcal{L}^{\infty}_{\sigma^{\ell_{0}}}$ and $A_{2}:=\operatorname{ess}\sup_{y\in\mathbb{R}^{d}}|f(y)|e^{-\tau^{\ell^{\prime}}(-y)}\leq\operatorname{ess}\sup_{y\in\mathbb{R}^{d}}|f(y)|<+\infty$ since $f\in\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ (by assumption). ∎
4.2. Alternative proofs for the Roumieu-type
We present now different techniques and ideas for the Roumieu case to show $(ii)\Rightarrow(i)$ in Theorem 4.2. Here we are not using functional analytic tools but involve special (optimal) functions. One has to work with different assumptions for the weight function matrices but the new introduced optimal functions can be used in other related contexts as well.
Let $\mathcal{W}=\{\omega^{\ell}:\ell>0\}$ be a weight function matrix. Then set
(4.7) $\theta^{p}_{\omega^{\ell}}(x):=e^{-p^{-1}\omega^{\ell}(x)},\;\;\;x\in\mathbb{R}^{d},\;\ell>0,\;1\leq p<\infty,$
and
(4.8) $\theta^{\infty}_{\omega^{\ell}}(x):=e^{-\omega^{\ell}(x)},\;\;\;x\in\mathbb{R}^{d},\;\ell>0.$
The extreme case $\omega^{\ell}=0$ gives $\theta^{p}_{\omega^{\ell}}=1=\theta^{\infty}_{\omega^{\ell}}$. In order to proceed, we require the following relations between the functions $\omega^{\ell}$:
(4.9) $\exists\;\ell^{\prime}_{0}>\ell_{0}>0:\;\;\;\sup_{x\in\mathbb{R}^{d}}\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)=+\infty,$
(4.10) $\exists\;\ell^{\prime}_{0}>\ell_{0}>0:\;\;\;\lim_{|x|\rightarrow+\infty}\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)=+\infty,$
and
(4.11) $\exists\;a>1\;\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;b\in\mathbb{R}\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\omega^{\ell}(x)-\omega^{\ell^{\prime}}(x)\geq a\log(1+|x|)+b.$
Remark 4.7.
We point out:
    If (4.9), (4.10) holds for some indices $\ell^{\prime}_{0}>\ell_{0}$, then also for all $\ell^{\prime},\ell$ satisfying $\ell^{\prime}\geq\ell^{\prime}_{0}>\ell_{0}\geq\ell>0$.
    By iteration it is clear that (4.11) is equivalent to
    (4.12) $\forall\;a>1\;\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;b\in\mathbb{R}\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\omega^{\ell}(x)-\omega^{\ell^{\prime}}(x)\geq a\log(1+|x|)+b.$
    Note that in (4.11), (4.12) necessarily the indices have to satisfy the relation $\ell>\ell^{\prime}$.
    One shall note that (4.9), (4.10) and (4.11) are crucial for the techniques in this section but they can never be satisfied for the simple weight function matrix $\mathcal{W}:=\{\omega\}$. On the other hand, the estimates in (3.13) and (3.14) hold for any sub-additive $\omega$ and hence the main results Theorems 4.1, 4.2 and Proposition 3.6 can be applied to the simple case, too.
Lemma 4.8.
Let $\mathcal{W}=\{\omega^{\ell}:\ell>0\}$ be a weight function matrix.
  • $(i)$
    We have
    $\forall\;\ell\geq\ell^{\prime}>0:\;\;\;\theta^{\infty}_{\omega^{\ell}}\in\mathcal{L}^{\infty}_{\omega^{\ell^{\prime}}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{W}\}}.$
  • $(ii)$
    If $\mathcal{W}$ satisfies (4.9) with indices $\ell^{\prime}_{0}>\ell_{0}$, then
    $\forall\;\ell^{\prime}\geq\ell^{\prime}_{0}>\ell_{0}\geq\ell>0:\;\;\;\theta^{\infty}_{\omega^{\ell}}\notin\mathcal{L}^{\infty}_{\omega^{\ell^{\prime}}}\supseteq\mathcal{L}^{\infty}_{(\mathcal{W})}.$
  • $(iii)$
    Assume that $\mathcal{W}$ is continuous and (4.11) holds. Then for any $1\leq p<\infty$ we get:
    $\forall\;\ell>0\;\exists\;\ell^{\prime}>0:\;\;\;\theta^{p}_{\omega^{\ell}}\in\mathcal{L}^{p}_{\omega^{\ell^{\prime}}}\subseteq\mathcal{L}^{p}_{\{\mathcal{W}\}}.$
  • $(iv)$
    Assume that $\mathcal{W}$ is continuous and (4.10) holds. Then for any $1\leq p<\infty$ we get:
    $\exists\;\ell_{0}^{\prime}>\ell_{0}\;\forall\;\ell^{\prime}\geq\ell^{\prime}_{0}>\ell_{0}\geq\ell>0:\;\;\;\theta^{p}_{\omega^{\ell}}\notin\mathcal{L}^{p}_{\omega^{\ell^{\prime}}}\supseteq\mathcal{L}^{p}_{(\mathcal{W})}.$
Proof. $(i)$ This is clear since $\sup_{x\in\mathbb{R}^{d}}|\theta^{\infty}_{\omega^{\ell}}(x)|e^{\omega^{\ell^{\prime}}(x)}=\sup_{x\in\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}}(x)-\omega^{\ell}(x)}\leq 1$.
$(ii)$ According to the given indices and the assumption it holds that $\sup_{x\in\mathbb{R}^{d}}|\theta^{\infty}_{\omega^{\ell}}(x)|e^{\omega^{\ell^{\prime}}(x)}=\sup_{x\in\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}}(x)-\omega^{\ell}(x)}\geq\sup_{x\in\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)}=+\infty$; see Remark 4.7.
$(iii)$ We apply (4.12) to (some fixed) $a\geq 2(d+1)$ and hence get for each $\ell>0$ some index $\ell^{\prime}>0$ such that
$\|\theta^{p}_{\omega^{\ell}}\|^{p}_{p,\omega^{\ell^{\prime}}}=\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}}(x)-\omega^{\ell}(x)}dx\leq e^{-b}\int_{\mathbb{R}^{d}}\frac{1}{(1+|x|)^{a}}dx.$
The last integral is finite by the choice $a\geq 2(d+1)$.
$(iv)$ Let $\ell^{\prime}\geq\ell^{\prime}_{0}>\ell_{0}\geq\ell>0$ be given, with $\ell^{\prime}_{0}$ and $\ell_{0}$ denoting the parameters appearing in (4.10) and then
$\|\theta^{p}_{\omega^{\ell}}\|^{p}_{p,\omega^{\ell^{\prime}}}=\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}}(x)-\omega^{\ell}(x)}dx\geq\int_{\mathbb{R}^{d}}e^{\omega^{\ell^{\prime}_{0}}(x)-\omega^{\ell_{0}}(x)}dx=+\infty.$
(However, in general it is not clear if already (4.9) implies this estimate.) ∎
Note:
    $(i)$ and $(iii)$ in Lemma 4.8 yield, in particular, the non-triviality of the corresponding weighted Roumieu-type spaces even in the anisotropic setting. This should be compared with Lemma 3.5: There it has been assumed that $\mathcal{W}$ is non-decreasing, whereas $(i)$ in Lemma 4.8 requires no further assumption on $\mathcal{W}$ and $(iii)$ there requires information on the difference between the growth of the weight functions which fails in the simple case; recall Remark 4.7.
    By $(ii)$ and $(iv)$ in Lemma 4.8 we see that in general $\theta^{\infty}_{\omega^{\ell}}$ resp. $\theta^{p}_{\omega^{\ell}}$ does not belong to the corresponding Beurling-type spaces and so the proofs of the next results are becoming unclear in this setting.
Using this preparation we prove the following variant of $(ii)$ in Proposition 4.6.
Proposition 4.9.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices.
Then the inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ (as sets) implies $\mathcal{S}\{\preceq\}\mathcal{T}$.
Proof. By $(i)$ in Lemma 4.8 and $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ we get $\theta^{\infty}_{\sigma^{\ell}}\in\mathcal{L}^{\infty}_{\{\mathcal{T}\}}$ for any $\ell>0$. Hence
$\forall\;\ell>0\;\exists\;\ell^{\prime}>0:\;\;\;\sup_{x\in\mathbb{R}^{d}}e^{\tau^{\ell^{\prime}}(x)-\sigma^{\ell}(x)}<+\infty,$
and so
$\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;D\geq 0\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell^{\prime}}(x)\leq\sigma^{\ell}(x)+D,$
which is precisely $\mathcal{S}\{\preceq\}\mathcal{T}$. ∎
Note: This statement is more general than $(ii)$ in Proposition 4.6 since no stability under translation is required.
Moreover, the established technique immediately gives the next result:
Proposition 4.10.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be weight function matrices. Then the following are equivalent:
  • $(i)$
    We have the continuous inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{(\mathcal{T})}$.
  • $(ii)$
    We have the inclusion $\mathcal{L}^{\infty}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{\infty}_{(\mathcal{T})}$ as sets.
  • $(iii)$
    The matrices satisfy
    (4.13) $\forall\;\ell>0\;\forall\;\ell^{\prime}>0\;\exists\;D\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{\ell^{\prime}}(x)+D;$
    and let us abbreviate this relation by $\mathcal{S}\vartriangleleft\mathcal{T}$.
Proof. $(iii)\Rightarrow(i)$ follows immediately by definition of the spaces and $(i)\Rightarrow(ii)$ is clear. For $(ii)\Rightarrow(iii)$ follow the proof of Proposition (4.9) and replace $\exists\;\ell^{\prime}>0$” by $\forall\;\ell^{\prime}>0$”. ∎
Finally, we treat the $L^{p}$-Roumieu-type spaces and so a variant of $(i)$ in Proposition 4.6. Here, due to technical reasons we focus on isotropic weight matrices and, before stating the next result, we introduce the following growth relations for $\mathcal{W}=\{\omega^{\ell}:\ell>0\}$:
(4.14) $\forall\;\ell>0\;\exists\;\ell_{1}>0\;\exists\;C\geq 1\;\forall\;t\in\mathbb{N}:\;\;\;\omega^{\ell_{1}}(t+1)\leq\omega^{\ell}(t)+C,$
(4.15) $\forall\;\ell_{1}>0\;\exists\;\ell>0\;\exists\;C\geq 1\;\forall\;t\in\mathbb{N}:\;\;\;\omega^{\ell_{1}}(t+1)\leq\omega^{\ell}(t)+C.$
    Similarly, as pointed out in Remark 3.8 (see (3.19) resp. (3.18)), one could consider the expression $C\omega^{\ell}(t)+C$ on the right-hand side. However, the additional appearance of the multiplicative constant $C$ is not appropriate in general; see again Section 5.1.
    (3.14) implies (4.14) and (3.13) implies (4.15) when we set $s:=1$.
    If $\mathcal{W}$ is non-decreasing, then equivalently in (4.14) and (4.15) we can replace $\forall\;t\in\mathbb{N}$” by $\forall\;t\geq 0$”; we only treat (4.14):
    When $t\geq 0$ with $t\notin\mathbb{N}$ is given we chose $s_{t}\in\mathbb{N}$ such that $s_{t}<t<s_{t}+1$ and get by iteration $\omega^{\ell_{2}}(s_{t}+1)\leq\omega^{\ell_{2}}(s_{t}+2)\leq\omega^{\ell_{1}}(s_{t}+1)+C\leq\omega^{\ell}(s_{t})+C_{1}+C\leq\omega^{\ell}(t)+C_{1}+C$. Hence it suffices to replace $C$ by $C_{1}+C$ and $\ell_{1}$ by $\ell_{2}(\leq\ell_{1})$ to ensure (4.14) for all $t\geq 0$.
Proposition 4.11.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be continuous and non-decreasing (and hence radial) weight function matrices. Moreover, assume that $\mathcal{S}$ satisfies (4.11) and that either $\mathcal{S}$ or $\mathcal{T}$ has (4.14).
Then the inclusion $\mathcal{L}^{p}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{p}_{\{\mathcal{T}\}}$ (as sets) for some/any $1\leq p<\infty$ implies $\mathcal{S}\{\preceq\}\mathcal{T}$.
Proof. By assumption (4.11), $(iii)$ in Lemma 4.8 and the inclusion $\mathcal{L}^{p}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{p}_{\{\mathcal{T}\}}$ (as sets) we obtain $\theta^{p}_{\sigma^{\ell}}\in\mathcal{L}^{p}_{\{\mathcal{T}\}}$ for all $\ell>0$ and so
(4.16) $\forall\;\ell>0\;\exists\;\ell_{1}>0:\;\;\;\int_{\mathbb{R}^{d}}|\theta^{p}_{\sigma^{\ell}}(x)|^{p}e^{\tau^{\ell_{1}}(x)}dx=\int_{\mathbb{R}^{d}}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}dx<+\infty.$
Let now $x\in\mathbb{R}^{d}$ with $n\leq|x|<n+1$ for some $n\in\mathbb{N}$. Then $e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}\geq e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}$ by using the assumption that both matrices are non-decreasing. Next recall that the volume of the $d$-dimensional sphere with radius $r$ is given by $V_{d}(r)=\frac{\pi^{d/2}r^{d}}{\Gamma(1+d/2)}$ (with $\Gamma$ denoting the Gamma-function) and therefore,
$\int_{x\in\mathbb{R}^{d}:n\leq|x|<n+1}1dx$ $=\int_{x\in\mathbb{R}^{d}:|x|<n+1}1dx-\int_{x\in\mathbb{R}^{d}:n\leq|x|}1dx=\frac{\pi^{d/2}((n+1)^{d}-n^{d})}{\Gamma(1+d/2)}$
$\geq\frac{\pi^{d/2}}{\Gamma(1+d/2)}=:C_{d},$
because $(n+1)^{d}-n^{d}\geq 1$ for all $n\in\mathbb{N}$ and $d\in\mathbb{N}_{>0}$. Then we estimate as follows:
$\int_{\mathbb{R}^{d}}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}dx$ $=\sum_{n=0}^{+\infty}\int_{x\in\mathbb{R}^{d}:n\leq|x|<n+1}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}dx\geq\sum_{n=0}^{+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}\int_{x\in\mathbb{R}^{d}:n\leq|x|<n+1}1dx$
$\geq C_{d}\sum_{n=0}^{+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}.$
Combining this estimate with (4.16) we get, in particular, that $\lim_{n\rightarrow+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}=0$ or equivalently $\lim_{n\rightarrow+\infty}\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)=$-$\infty$. Summarizing, so far we have shown
(4.17) $\forall\;\ell>0\;\exists\;\ell_{1}>0\;\exists\;D\geq 0\;\forall\;n\in\mathbb{N}:\;\;\;\tau^{\ell_{1}}(n)\leq\sigma^{\ell}(n+1)+D.$
Now assume that (4.14) for $\mathcal{S}$ holds and let $x\in\mathbb{R}^{d}$ with $n\leq|x|<n+1$ for some $n\in\mathbb{N}$. Then when iterating this property we estimate by
$\tau^{\ell_{1}}(x)$ $\leq\tau^{\ell_{1}}(n+1)\leq\sigma^{\ell}(n+2)+D\leq\sigma^{\ell^{\prime}}(n+1)+C+D\leq\sigma^{\ell^{\prime\prime}}(n)+C_{1}+C+D$
$\leq\sigma^{\ell^{\prime\prime}}(x)+C_{1}+C+D,$
thus $\mathcal{S}\{\preceq\}\mathcal{T}$ is verified (between the indices $\ell^{\prime\prime}$ and $\ell_{1}$).
If $\mathcal{T}$ satisfies (4.14), then similarly by (4.17) for any $x\in\mathbb{R}^{d}$ with $n\leq|x|<n+1$ for some $n\in\mathbb{N}_{>0}$ we have
$\tau^{\ell^{\prime\prime}_{1}}(x)$ $\leq\tau^{\ell^{\prime\prime}_{1}}(n+1)\leq\tau^{\ell^{\prime}_{1}}(n)+C\leq\tau^{\ell_{1}}(n-1)+C_{1}+C\leq\sigma^{\ell}(n)+C_{1}+C+D$
$\leq\sigma^{\ell}(x)+C_{1}+C+D.$
And for $x\in\mathbb{R}^{d}$ with $0\leq|x|<1$ one has $\tau^{\ell^{\prime\prime}_{1}}(x)\leq\tau^{\ell^{\prime\prime}_{1}}(1)\leq\sigma^{\ell}(x)+A$ with $A:=\tau^{\ell^{\prime\prime}_{1}}(1)$. Summarizing, again $\mathcal{S}\{\preceq\}\mathcal{T}$ is verified (between the indices $\ell$ and $\ell^{\prime\prime}_{1}$). ∎
Remark 4.12.
Without the additional assumptions on the weight matrices it is not clear that (4.16) implies $\lim_{|x|\rightarrow+\infty}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}=0$ or, what is enough to conclude, $\sup_{x\in\mathbb{R}^{d}}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}<+\infty$. This is due to the fact that in general $t\mapsto e^{\tau^{\ell_{1}}(t)-\sigma^{\ell}(t)}$ might behave very irregular: This function can have (infinitely many) “exploding peaks” but being sufficiently small “nearly everywhere on $\mathbb{R}$” and hence (4.16) still can be valid. In view of Remark 4.7 this comment becomes relevant in the simple case.
Proposition 4.13.
Let $\mathcal{S}=\{\sigma^{\ell}:\ell>0\}$ and $\mathcal{T}=\{\tau^{\ell}:\ell>0\}$ be continuous and non-decreasing (and hence radial) weight function matrices. Moreover, assume that $\mathcal{S}$ satisfies (4.11) and (4.14). Then the following are equivalent:
  • $(i)$
    We have the continuous inclusion $\mathcal{L}^{p}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{p}_{(\mathcal{T})}$ for some/any $1\leq p<\infty$.
  • $(ii)$
    We have the inclusion $\mathcal{L}^{p}_{\{\mathcal{S}\}}\subseteq\mathcal{L}^{p}_{(\mathcal{T})}$ as sets for some/any $1\leq p<\infty$.
  • $(iii)$
    The matrices satisfy $\mathcal{S}\vartriangleleft\mathcal{T}$.
The same equivalences are valid if $\mathcal{S}$ is assumed as above, but not necessarily having (4.14), and $\mathcal{T}$ satisfies (4.15) instead.
Proof. $(iii)\Rightarrow(i)$ follows by definition of the spaces and $(i)\Rightarrow(ii)$ is clear. For $(ii)\Rightarrow(iii)$ replace $\exists\;\ell_{1}>0$” by $\forall\;\ell_{1}>0$” in the proof of Proposition 4.11 and so get the stronger variant of (4.17). Then involve in the last step (4.14) for $\mathcal{S}$ resp. (4.15) for $\mathcal{T}$. ∎
5. Special cases of weight function matrices
We investigate two special but crucial situations for the defining (non-simple and non-stabilizing) weight function matrix $\mathcal{W}:=\{\omega^{\ell}:\ell>0\}$. Let $\omega:\mathbb{R}^{d}\rightarrow[0,+\infty)$ be a fixed function and consider
(5.1) $\omega^{\ell}(x):=\ell\omega(x),\;\;\;\ell>0,\;x\in\mathbb{R}^{d},$
resp.
(5.2) $\omega^{\ell}(x):=\omega(\ell x),\;\;\;\ell>0,\;x\in\mathbb{R}^{d}.$
If (5.1) is valid, then $\mathcal{W}$ is called to be of exponential-type and write $\mathcal{W}^{\mathfrak{c}}_{\omega}$ instead of $\mathcal{W}$. If (5.2) is valid, then $\mathcal{W}$ is called to be of dilatation-type and we write $\mathcal{W}_{\omega,\mathfrak{c}}$.
This terminology is inspired by the notation introduced in [31] and [32] and by taking into account the fact that the crucial seminorms are expressed by the function $\exp\circ\omega^{\ell}$. Note that known examples in the literature are given in terms of an exponential-type (and isotropic) weight matrix; see Section 6.3 for more details.
5.1. On the exponential-type case
$\mathcal{W}^{\mathfrak{c}}_{\omega}$ is a weight function matrix according to Definition 3.1. We gather some immediate facts and conclusions:
  • $(a)$
    $\mathcal{W}^{\mathfrak{c}}_{\omega}$ is non-decreasing/continuous/radial if and only if $\omega$ is so.
  • $(b)$
    The following are equivalent:
      $\sup_{x\in\mathbb{R}^{d}}\omega(x)<+\infty$,
      (3.4) holds,
      (3.5) holds.
    Therefore, $\mathcal{W}^{\mathfrak{c}}_{\omega}$ is unbounded if and only if $\sup_{x\in\mathbb{R}^{d}}\omega(x)=+\infty$.
  • $(c)$
    Let $\omega$ be non-decreasing and radial, then the following are equivalent:
      $\omega$ satisfies $(\omega_{1})$.
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies
      $\exists\;n_{0},\ell_{0}>0\;\exists\;C\geq 1\;\forall\;t\geq 0:\;\;\;\omega^{\ell_{0}}(2t)\leq\omega^{n_{0}}(t)+C.$
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.15).
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.13).
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.16).
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.17).
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.14).
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.18).
      $\mathcal{W}^{\mathfrak{c}}_{\omega}$ satisfies (3.19).
    These equivalences follow by definition and in view of Lemma 3.7; recall $(f)$ in Section 2.4.
  • $(d)$
    The following are equivalent:
      $\mathcal{W}^{\mathfrak{c}}_{\sigma}(\preceq)\mathcal{W}^{\mathfrak{c}}_{\tau}$ holds, i.e. (4.1).
      $\mathcal{W}^{\mathfrak{c}}_{\sigma}\{\preceq\}\mathcal{W}^{\mathfrak{c}}_{\tau}$ holds, i.e. (4.2).
      The matrices $\mathcal{W}^{\mathfrak{c}}_{\sigma}$ and $\mathcal{W}^{\mathfrak{c}}_{\tau}$ are related by
      (5.3) $\exists\;\ell_{0}>0\;\exists\;n_{0}>0\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell_{0}}(x)\leq\sigma^{n_{0}}(x)+C.$
[…]
    Moreover, by analogous reasons, the following are equivalent:
      $\mathcal{W}^{\mathfrak{c}}_{\sigma}\vartriangleleft\mathcal{W}^{\mathfrak{c}}_{\tau}$,
      $\exists\;n_{0}>0\;\forall\;\ell>0\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{n_{0}}(x)+C,$
      $\exists\;\ell_{0}>0\;\forall\;n>0\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell_{0}}(x)\leq\sigma^{n}(x)+C,$
[…]
  • $(e)$
    (4.9) holds if and only if $\sup_{x\in\mathbb{R}}\omega(x)=+\infty$ and (4.10) if and only if $\lim_{|x|\rightarrow+\infty}\omega(x)=+\infty$.
  • $(f)$
    $(\omega_{3,w})$ for $\omega$ in the anisotropic setting precisely means
    (5.4) $\exists\;a>0\;\exists\;b\in\mathbb{R}\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\omega(x)\geq a\log(1+|x|)+b,$
    and $(\omega_{3})$ is
    (5.5) $\forall\;a\geq 1\;\exists\;b\in\mathbb{R}\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\omega(x)\geq a\log(1+|x|)+b.$
    We verify that (4.12), which is a reformulation of (4.11), is equivalent to (5.5): On the one hand, when $a\geq 1$ is given and (4.12) holds then apply this condition to $a^{\prime}>1$ (large) and $\ell>\ell^{\prime}>0$ (both sufficiently small) such that $\frac{a^{\prime}}{\ell-\ell^{\prime}}\geq a$ is valid and which verifies (5.5) for this particular $a$. Conversely, if (5.5) holds and if $0<\ell^{\prime}<\ell$ (both small) and $a\geq 1$ are given, then there exists $a^{\prime}>1$ such that $a^{\prime}\geq\frac{a}{\ell-\ell^{\prime}}$ and when applying (5.5) to $a^{\prime}$ property (4.12) follows with these choices for $a,\ell,\ell^{\prime}$.
  • $(g)$
    The following are equivalent for radial $\omega$:
      (4.14) for $\mathcal{W}^{\mathfrak{c}}_{\omega}$ is valid.
      (4.15) for $\mathcal{W}^{\mathfrak{c}}_{\omega}$ is valid.
      $\omega$ satisfies
      (5.6) $\exists\;C\geq 1\;\forall\;n\in\mathbb{N}:\;\;\;\omega(n+1)\leq C\omega(n)+C.$
    If $\omega$ is non-decreasing and has $(\omega_{1})$, then (5.6) follows but the latter condition is much weaker: For example $\omega(t):=e^{t}$ clearly has (5.6) with $C:=e$ but $(\omega_{1})$ fails. And, analogously as in Section 4.2, in (5.6) equivalently we can replace $n$ by any $t\geq 0$ when modifying $C$.
Summarizing, the main statements Theorems 4.1 and 4.2 can be applied to exponential-type weight matrices and the characterizing relation of the inclusions of the weighted spaces is given by $\preceq$ between the defining weight functions. In the radial setting property $(\omega_{1})$ is crucial to ensure stability under translation.
5.2. On the dilatation-type case
We give now the analogous comments for the weight function matrix $\mathcal{W}_{\omega,\mathfrak{c}}$. First note that $\mathcal{W}_{\omega,\mathfrak{c}}$ is formally a weight matrix according to Definition 3.1 if $\omega$ is non-decreasing. Moreover, we get:
  • $(a)$
    $\mathcal{W}_{\omega,\mathfrak{c}}$ is non-decreasing/continuous/radial if and only if $\omega$ is so.
  • $(b)$
    The following are equivalent:
      $\sup_{x\in\mathbb{R}^{d}}\omega(x)<+\infty$,
      (3.4) holds,
      (3.5) holds.
    Therefore, $\mathcal{W}_{\omega,{\mathfrak{c}}}$ is unbounded if and only if $\sup_{x\in\mathbb{R}^{d}}\omega(x)=+\infty$.
  • $(c)$
    Let $\omega$ be non-decreasing and radial, then $\mathcal{W}_{\omega,\mathfrak{c}}$ satisfies (3.15) and(3.17) since both conditions amount to have $\omega(2\ell t)\leq\omega(nt)+L$ and which is clear: choose $n\geq 2\ell$ when $\ell>0$ is given and $\ell\leq\frac{n}{2}$ when $n>0$ is given. (Indeed, $L\geq 1$ is superfluous and can be taken arbitrarily.)
    Thus, also (3.18) and (3.19) follow.
  • $(d)$
[…]
  • $(e)$
    (4.9) holds if and only if $\sup_{x\in\mathbb{R}}\omega(ax)-\omega(x)=+\infty$ for some $a>1$ and (4.10) if and only if $\lim_{|x|\rightarrow+\infty}\omega(ax)-\omega(x)=+\infty$ for some $a>1$.
    In general, both requirements are not clear but when $\omega\equiv\omega_{\mathbf{M}}$, $\omega_{\mathbf{M}}$ denoting the associated weight function, see (7.22), and with $\mathbf{M}\in\mathbb{R}_{>0}^{\mathbb{N}}$ being a log-convex weight sequence in the notion of [31, Def. 2.4], then both conditions are valid by recalling (the proof of) [31, Lemma 2.9]. Indeed, in this case even any choice $a>1$ is sufficient to conclude.
  • $(f)$
    $\omega$ satisfies (5.4) if and only if $\omega^{\ell}$ does so for some/any $\ell>0$ and similarly for (5.5). However, the equivalence to (4.12) established in $(f)$ in the previous Section 5.1 is not clear in general even if $\omega\equiv\omega_{\mathbf{M}}$ but in this situation the following is known: (4.12) holds if and only if $\mathbf{M}$ is derivation closed which is the classical condition $(M.2)^{\prime}$ in [15]; see [4, Lemma 2] even in a more general (mixed) situation.
  • $(g)$
    If $\omega$ is non-decreasing and radial, then (4.14) resp. (4.15) for $\mathcal{W}^{\mathfrak{c}}_{\omega}$ amounts to have $\omega(\ell_{1}(n+1))\leq\omega(\ell n)+C$ for some $C\geq 1$ and all $n\in\mathbb{N}$. Thus, both properties hold automatically when $\omega$ is assumed to be non-decreasing: In the first estimate choose $\ell_{1}:=\frac{\ell}{2}$ and in the second one $\ell:=2\ell_{1}$.
Summarizing, the main statements Theorems 4.1 and 4.2 can be applied to dilatation-type weight matrices; here the inclusion is characterized in terms of relation $\preceq_{\mathfrak{c}}$ between the defining weight functions. In view of comment $(d)$ under additional assumptions on the weights a connection to the corresponding exponential-type weight system and the (characterizing) relation $\preceq$ can be established. And by comment $(c)$ it holds that the associated weighted classes are invariant under translation for any non-decreasing and radial $\omega$.
6. Weighted (test-)function classes of Beurling-Björck-type
We study related and more general cases and comment how and if the results from Sections 3 and 4 can be transferred.
6.1. Weighted subspaces of a fixed Banach space
Let $\mathcal{W}=\{\omega^{\ell}:\ell>0\}$ be a weight function matrix and let $\mathcal{A}^{p}\subseteq L^{p}$ be a fixed Banach space, $1\leq p\leq\infty$. Then define
$\mathcal{A}^{\infty}_{\omega^{\ell}}:=\left\{f\in\mathcal{A}^{\infty}:\;\;\;\|f\|_{\infty,\omega^{\ell}}<+\infty\right\},$
and
$\mathcal{A}^{\infty}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{A}^{\infty}_{\omega^{\ell}},\hskip 15.0pt\mathcal{A}^{\infty}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{A}^{\infty}_{\omega^{\ell}},$
endowed with their natural topologies. Moreover, if $\mathcal{W}$ is continuous and $1\leq p<\infty$, then for any parameter $\ell>0$ let us set
$\mathcal{A}^{p}_{\omega^{\ell}}:=\left\{f\in\mathcal{A}^{p}:\;\;\;\|f\|_{p,\omega^{\ell}}<+\infty\right\},$
and put accordingly
$\mathcal{A}^{p}_{\{\mathcal{W}\}}:=\bigcup_{\ell>0}\mathcal{A}^{p}_{\omega^{\ell}},\hskip 15.0pt\mathcal{A}^{p}_{(\mathcal{W})}:=\bigcap_{\ell>0}\mathcal{A}^{p}_{\omega^{\ell}}.$
Again, $\mathcal{A}^{p}_{\{\mathcal{W}\}}$ and $\mathcal{A}^{p}_{(\mathcal{W})}$ are endowed with their natural topologies and the following continuous inclusions hold:
$\mathcal{A}^{p}_{(\mathcal{W})}\subseteq\mathcal{A}^{p}_{\{\mathcal{W}\}}\subseteq\mathcal{A}^{p},\;\;\;1\leq p\leq\infty.$
Thus, the special case $\mathcal{A}^{p}=L^{p}$ corresponds to the situation studied in Sections 3 and 4 and obviously, by definition,
(6.1) $\mathcal{A}^{p}_{[\mathcal{W}]}=\mathcal{L}^{p}_{[\mathcal{W}]}\cap\mathcal{A}^{p}\subseteq\mathcal{L}^{p}_{[\mathcal{W}]},\;\;\;1\leq p\leq\infty.$
We comment in detail how the shown statements transfer if the symbol/functor $\mathcal{L}$ is replaced by $\mathcal{A}$. A natural basic requirement, to avoid trivial situations and which is assumed from now on, is $\mathcal{A}^{p}\neq\{0\}$; i.e. the fixed space is non-trivial. Otherwise, of course $\mathcal{A}^{p}_{[\mathcal{W}]}=\{0\}$ holds.
  • $(a)$
    Section 3.2; Theorem 3.3: By definition, (3.4) implies $\mathcal{A}^{p}_{\{\mathcal{W}\}}=\mathcal{A}^{p}$ (as sets and l.c.v.s.) and (3.5) yields the same equalities for the Beurling-types.
    Concerning the converse, the proof of $(i)$ in Proposition 3.2 immediately transfers when we assume that $\mathcal{A}^{\infty}$ contains at least all (equivalence classes of) constant functions and, indeed, it suffices to assume
    (6.2) $\exists\;(0\neq)\widetilde{f}\in\mathcal{C}(\mathbb{R}^{d}):\;\;\;\widetilde{f}\in\mathcal{A}^{\infty},\;\;\;\inf_{x\in\mathbb{R}^{d}}|\widetilde{f}(x)|>\epsilon>0.$
    Note that even in the proof of $(i)$ in Proposition 3.2 it is not sufficient to take any arbitrary (non-trivial) $f\in\mathcal{C}(\mathbb{R}^{d})\cap L^{\infty}$ since such $f$ can vanish identically near infinity. To transfer the proof of $(ii)$ we assume that $\mathcal{A}^{p}$ contains at least all (equivalence classes of) locally constant functions $f$ satisfying $\|f\|_{L^{p}}<+\infty$. Thus $g^{p}\in\mathcal{A}^{p}$, $g^{p}$ the function defined in (3.7). Therefore, in this case the characterizing Theorem 3.3 is valid, too.
  • $(b)$
    Section 3.3: The non-triviality of $\mathcal{A}^{p}_{[\mathcal{W}]}$ is not clear in general even if $\mathcal{A}^{p}\neq\{0\}$ because the technical constructive proof of Lemma 3.5 cannot be transferred automatically.
    In view of (6.1) the non-triviality amounts to having
    $\{0\}\neq\mathcal{L}^{p}_{[\mathcal{W}]}\cap\mathcal{A}^{p}.$
  • $(c)$
    Section 3.4: By the analogous definition and system of seminorms and by recalling that the proof of Proposition 3.6 purely involves weights and their growth properties, it follows that this result applies to $\mathcal{A}^{p}$.
  • $(d)$
    Section 4.1; main Theorems 4.1 and 4.2: Since the spaces are expressed by the analogous system of seminorms and by taking into account the previous comment $(c)$ we have that the crucial results Propositions 4.3, 4.5 and 4.6 hold in this setting, too. Note (again) that the proofs of Propositions 4.5 and 4.6 require the non-triviality of the corresponding (smaller) weighted classes and which is not clear by comment $(b)$ above. However, when taking into account the conventions mentioned in Remark 4.4 one is able to conclude in any case.
  • $(e)$
    Section 4.2: Similarly as in comment $(b)$ before in general the methods involving the optimal functions $\theta^{p}_{\omega^{\ell}}$ from (4.7) and $\theta^{\infty}_{\omega^{\ell}}$ from (4.8) fail since neither $\theta^{p}_{\omega^{\ell}}\in\mathcal{A}^{p}_{\{\mathcal{W}\}}$ nor $\theta^{\infty}_{\omega^{\ell}}\in\mathcal{A}^{\infty}_{\{\mathcal{W}\}}$ is clear. If these functions belong to the spaces, then all shown results transfer immediately.
    On the other hand, when one assumes $\widetilde{\theta}^{\infty}_{\omega^{\ell}}:=\widetilde{f}\cdot\theta^{\infty}_{\omega^{\ell}}\in\mathcal{A}^{\infty}_{\{\mathcal{W}\}}$, $\widetilde{f}\in\mathcal{C}(\mathbb{R}^{d})$ a function satisfying the infimum-condition from (6.2), then the crucial estimate in the proof of Proposition 4.9 reads as follows:
    $\forall\;\ell>0\;\exists\;\ell^{\prime}>0\;\exists\;D\geq 0\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell^{\prime}}(x)\leq\sigma^{\ell}(x)-\log(|\widetilde{f}(x)|)+D\leq\sigma^{\ell}(x)-\log(\epsilon)+D,$
    which yields $\mathcal{S}\{\preceq\}\mathcal{T}$. Note that $\log(\epsilon)>0$ because w.l.o.g. $\epsilon\in(0,1)$. Similarly Proposition 4.10 holds, too.
    And assume that there exists a continuous function $(0\neq)\widetilde{f}\in\mathcal{C}(\mathbb{R}^{d})$ with $\inf_{x\in\mathbb{R}^{d}}|\widetilde{f}(x)|>\epsilon^{1/p}$ for some $\epsilon>0$ and such that $\widetilde{\theta}^{p}_{\omega^{\ell}}:=\widetilde{f}\cdot\theta^{p}_{\omega^{\ell}}\in\mathcal{A}^{p}_{\{\mathcal{W}\}}$ for all $\ell>0$. Then also Propositions 4.11 and 4.13 are valid (under the same assumptions on $\mathcal{S}$ and $\mathcal{T}$): First we have that (4.16) turns into
    (6.3) $\forall\;\ell>0\;\exists\;\ell_{1}>0:\;\;\;\int_{\mathbb{R}^{d}}|\widetilde{\theta}^{p}_{\sigma^{\ell}}|^{p}dx=\int_{\mathbb{R}^{d}}|\widetilde{f}(x)|^{p}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}dx<+\infty.$
    Second, the crucial estimate in the proof of Proposition 4.11 changes as follows:
    $\int_{\mathbb{R}^{d}}|\widetilde{f}(x)|^{p}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}dx=\sum_{n=0}^{+\infty}\int_{x\in\mathbb{R}^{d}:n\leq|x|<n+1}|\widetilde{f}(x)|^{p}e^{\tau^{\ell_{1}}(x)-\sigma^{\ell}(x)}dx$
    $\geq\sum_{n=0}^{+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}\int_{x\in\mathbb{R}^{d}:n\leq|x|<n+1}|\widetilde{f}(x)|^{p}dx$
    $\geq\sum_{n=0}^{+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}\inf_{x\in\mathbb{R}^{d}}|\widetilde{f}(x)|^{p}\int_{x\in\mathbb{R}^{d}:n\leq|x|<n+1}1dx$
    $=\sum_{n=0}^{+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}\inf_{x\in\mathbb{R}^{d}}|\widetilde{f}(x)|^{p}\frac{\pi^{d/2}((n+1)^{d}-n^{d})}{\Gamma(1+d/2)}\geq\sum_{n=0}^{+\infty}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}\frac{\pi^{d/2}\epsilon}{\Gamma(1+d/2)}.$
    For the last estimate recall that $(n+1)^{d}-n^{d}\geq 1$ for all $n\in\mathbb{N}$ and $d\in\mathbb{N}_{>0}$.
    Combining this estimate with (6.3) it follows again that $\sup_{n\in\mathbb{N}}e^{\tau^{\ell_{1}}(n)-\sigma^{\ell}(n+1)}<+\infty$; i.e. (4.17) is verified and the rest follows as in the proof above.
6.2. Weighted Fourier image spaces
In this section we comment on weighted spaces which are defined analogously as the ones in Section 3 but the crucial difference is that we are weighting the Fourier transform $\widehat{f}$ of the functions under consideration:
For a fixed function space $\mathcal{A}$, a (continuous) weight function matrix $\mathcal{W}$ and $1\leq p\leq\infty$ we set
[…]
and both spaces are endowed with their natural topologies. Of course, in order to make this definition useful one requires that the Fourier-transform is well-defined on $\mathcal{A}$ and that $\widehat{\mathcal{A}}^{p}_{[\mathcal{W}]}$ is non-trivial. We remark:
  • $(a)$
    Proposition 4.3 is clear for this setting by the analogous definition of the seminorms but with $f$ being replaced by $\widehat{f}$.
  • $(b)$
    The non-triviality of $\widehat{\mathcal{A}}^{p}_{[\mathcal{W}]}$ is not guaranteed in general and has to be checked. This step requires technical investigations, in particular, when explicitly constructing (optimal) functions belonging to such classes.
  • $(c)$
    Assume that $\{0\}\neq\widehat{\mathcal{A}}^{p}_{[\mathcal{S}]}$ holds, then the proofs of Propositions 4.5 and 4.6 can be transferred as follows:
    Take a (fixed) non-trivial $f\in\widehat{\mathcal{A}}^{p}_{[\mathcal{S}]}$ and for any $x_{0}\in\mathbb{R}^{d}$ put $g_{x_{0}}(x):=e^{i\langle x,x_{0}\rangle}f(x)$. Then $\widehat{g}_{x_{0}}(x)=\widehat{f}(x-x_{0})$ and apply (4.3), (4.4), (4.5), and (4.6) to the family $\{g_{x_{0}}:x_{0}\in\mathbb{R}^{d}\}$; see also the original proof of [3, Thm. 1.3.18] (for the Beurling-type).
    However, as mentioned in Remark 4.4 and in comment $(d)$ in Section 6.1, the same implication holds when $\{0\}=\widehat{\mathcal{A}}^{p}_{[\mathcal{S}]}$ (and $\{0\}=\widehat{\mathcal{A}}^{p}_{[\mathcal{T}]}$).
    Summarizing, the analogues of the main results Theorems 4.1 and 4.2 hold for this described “Fourier-setting” too when assuming the same conditions on the matrices $\mathcal{S}$ and $\mathcal{T}$.
    Indeed, Proposition 3.6 transfers and shows that under (3.13) resp. (3.14) the corresponding weighted classes are modulation invariant.
6.3. Testfunction spaces appearing in the literature
We revisit now the different types and notions of prominent weighted (test-)function spaces which have been considered in the literature. All known examples are expressed in terms of radial/isotropic exponential-type weight function matrices and by the Fourier framework described in the previous Section 6.2 when taking spaces $\mathcal{A}\subseteq L^{1}$. Finally, $(A)$ and $(B)$ correspond to $p=1$ and $(C)$ to $p=\infty$.
  • $(A)$
    Let $\omega$ be a non-quasianalytic BMT-weight except necessarily satisfying the convexity condition $(\omega_{4})$. Let $U\subseteq\mathbb{R}^{d}$ be non-empty open and then set
    $\mathcal{D}_{\{\omega\}}(U):=\{f\in\mathcal{D}(U),\;\,\exists\;\ell>0:\;\int_{\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}dx<+\infty\},$
    and
    $\mathcal{D}_{(\omega)}(U):=\{f\in\mathcal{D}(U),\;\;\forall\;\ell>0:\;\int_{\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}dx<+\infty\}.$
    For this setting see [7, Def. 3.1] and in this work, as described in the introduction, with the help of $(\omega_{4})$ those spaces have been transferred into the nowadays frequently used definition involving $\varphi^{*}_{\omega}$; they are usually denoted by $\mathcal{E}_{\{\omega\}}(U)$, $\mathcal{E}_{(\omega)}(U)$ (also for quasianalytic weights).
  • $(B)$
    In [3, Def. 1.3.1, 1.3.2 & 1.3.4] for any $\omega:[0,+\infty)\rightarrow[0,+\infty)$ which satisfies $(\omega_{\operatorname{sub}})$, i.e. $(\alpha)$ in [3], the following (global) test function classes (of Beurling-type) have been introduced:
    $(\mathfrak{D}_{(\omega)}(U)=)\mathfrak{D}_{\omega}(U):=\{f\in L^{1}(\mathbb{R}^{d}):\exists\;K\subset\subset U,\;\operatorname{supp}(f)\subseteq K,\;\forall\;\ell>0:\;\int_{\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}dx<+\infty\}.$
    The non-triviality of this class has been characterized in [3, Thm. 1.3.7] in terms of the non-quasianalyticity condition for weight functions.
  • $(C)$
    Finally, in [23] for $\omega$ satisfying conditions (1)(4)(1)-(4) from Section 2.8 the following test function classes have been studied for $U\subseteq\mathbb{R}^{d}$ (see [23, p. 18-19]):
    $\widetilde{\mathcal{D}}_{\{\omega\}}(U):=\{f\in L^{1}(\mathbb{R}^{d}):\exists\;K\subset\subset U,\;\operatorname{supp}(f)\subseteq K,\;\exists\;\ell>0:\;\sup_{x\in\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}<+\infty\},$
    and
    $\widetilde{\mathcal{D}}_{(\omega)}(U):=\{f\in L^{1}(\mathbb{R}^{d}):\exists\;K\subset\subset U,\;\operatorname{supp}(f)\subseteq K,\;\forall\;\ell>0:\;\sup_{x\in\mathbb{R}^{d}}|\widehat{f}(x)|e^{\ell\omega(x)}<+\infty\}.$
Consequently, the established results can be applied to all these examples and since we deal with exponential-type weight matrices relation $\preceq$ between the defining weights is characterizing the inclusion(s) of the above weighted spaces and equivalence (as l.c.v.s.) is characterized in terms of equivalent weight functions, i.e. relation $\sim$ is relevant; see the comments in Sections 5.1 and 6.2.
Some parts of these characterizations have already been stated in the literature, for $(B)$ we refer to [3, Thm. 1.3.18] concerning the Beurling-type, for $(A)$ in [7, Rem. 3.2 $(2)$] it has been mentioned that U. Franken has also treated the Beurling-type in his Master thesis [9]. Note that non-quasianalyticity $(\omega_{\operatorname{nq}})$ for the defining weight $\omega$ is crucial in each setting in order to ensure non-triviality.
Remark 6.1.
Indeed, in [23] the authors have even considered the non-standard situation when the (matrix) parameter $\ell$ is restricted to $(0,\Lambda)$ for $\widetilde{\mathcal{D}}_{(\omega)}(U)$ and to $(\Lambda,+\infty)$ for $\widetilde{\mathcal{D}}_{\{\omega\}}(U)$. Here, $\Lambda\in(0+\infty)$ denotes a “truncation parameter”.
The established techniques transfer immediately with the obvious changes in the proofs of Proposition 3.6, Lemma 3.7, Theorems 4.1 and 4.2 when restricting in the particular conditions the indices w.r.t. given truncation parameters $\Lambda^{\prime}$ and $\Lambda$ for the matrices $\mathcal{T}$ and $\mathcal{S}$, respectively, and so also the above comments in Section 6 apply. More precisely, the characterizing growth relation $\mathcal{S}(\preceq)\mathcal{T}$ and hence (4.1) turns into
(6.4) $\forall\;\ell\in(0,\Lambda^{\prime})\;\exists\;n\in(0,\Lambda)\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{n}(x)+C,$
whereas (4.2) turns into
(6.5) $\forall\;n\in(\Lambda,+\infty)\;\exists\;\ell\in(\Lambda^{\prime},+\infty)\;\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{n}(x)+C.$
Now focus on the exponential-type system and first note that the equivalences stated in $(c)$ and $(d)$ in Section 5.1 are not valid anymore in this situation: Recall that $(\omega_{1})$ for a radial $\omega$ means $\omega(2t)\leq L\omega(t)+L$ for some $L\geq 1$ and all $t\geq 0$. This constant $L$ yields then a uniform scaling in the parameters of the matrix $\mathcal{W}^{\mathfrak{c}}_{\omega}$; i.e. one infers the relation $L\ell=n$ and which is a problem when restricting the parameters and when $\ell\rightarrow\Lambda^{\prime}$ (resp. $n\rightarrow\Lambda$). Indeed, $L$ yields a transformation (dilation) of the truncation parameter.
[…]
  • $(i)$
    One has
    (6.6) $\exists\;C\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau(x)\leq\frac{\Lambda}{\Lambda^{\prime}}\sigma(x)+C.$
  • $(ii)$
    (6.4) holds.
  • $(iii)$
    (6.5) holds.
  • $(iv)$
    One has
    (6.7)  $\forall\;0<K^{\prime}<\Lambda^{\prime}\;\forall\;K>\Lambda\;\exists\;D\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau(x)\leq\frac{K}{K^{\prime}}\sigma(x)+D.$
Then the implications $(i)\Rightarrow(ii),(iii)\Rightarrow(iv)$ are valid:
If (6.6) holds, then (6.4) follows by taking $n:=\frac{\Lambda}{\Lambda^{\prime}}\ell\in(0,\Lambda)$ for any given $\ell\in(0,\Lambda^{\prime})$ and (6.5) follows when taking $\ell:=n\frac{\Lambda^{\prime}}{\Lambda}\in(\Lambda^{\prime},+\infty)$ for any given $n\in(\Lambda,+\infty)$.
$(ii),(iii)\Rightarrow(iv)$: Fix $K^{\prime}<\Lambda^{\prime}$ and $K>\Lambda$ and then by (6.4) one has with $\ell:=\Lambda^{\prime}-\epsilon_{1}$ and $n:=\Lambda-\epsilon_{2}$, both $\epsilon_{i}$ chosen sufficiently small depending on given $K^{\prime}$, that
$\exists\;\epsilon_{1},\epsilon_{2}>0:\;\;\;\frac{\Lambda-\epsilon_{2}}{\Lambda^{\prime}-\epsilon_{1}}\leq\frac{K}{K^{\prime}}.$
Note that by the point-wise order of the weight functions in the matrices in (6.4) we are interested in $\ell\rightarrow\Lambda^{\prime}$ and one can assume w.l.o.g. $n=n(\ell)\rightarrow\Lambda$. Similarly, (6.5) implies with $n:=\Lambda+\epsilon_{1}$ and $\ell:=\Lambda^{\prime}+\epsilon_{2}$, again both $\epsilon_{i}$ small enough depending on given $K$, that
$\exists\;\epsilon_{1},\epsilon_{2}>0:\;\;\;\frac{\Lambda+\epsilon_{1}}{\Lambda^{\prime}+\epsilon_{2}}\leq\frac{K}{K^{\prime}}.$
Thus (6.7) holds in both cases.
Finally, (4.13) and so $\mathcal{S}\vartriangleleft\mathcal{T}$ transfers in this setting to
(6.8) $\forall\;\ell\in(0,\Lambda^{\prime})\;\forall\;n\in(\Lambda,+\infty)\;\exists\;D\geq 1\;\forall\;x\in\mathbb{R}^{d}:\;\;\;\tau^{\ell}(x)\leq\sigma^{n}(x)+D.$
Indeed (6.7) and (6.8) are equivalent: Let $\epsilon,\epsilon_{1}>0$, then when applying (6.7) to $K:=\Lambda+\epsilon$ and $K^{\prime}:=\Lambda^{\prime}-\epsilon_{1}$ we get (6.8) with $\ell=K^{\prime}$ and $n=K$. Conversely, we apply (6.8) to $\ell=\Lambda^{\prime}-\epsilon_{1}$ and $n=\Lambda-\epsilon$ and get (6.7) with $K^{\prime}=\ell$ and $K=n$. Since $\epsilon,\epsilon_{1}$ can be chosen arbitrarily small, the equivalence is established.
This also illustrates the difference compared with the standard setting when not involving a truncation parameter since in this case $\mathcal{S}\vartriangleleft\mathcal{T}$ obviously implies both $\mathcal{S}(\preceq)\mathcal{T}$ and $\mathcal{S}\{\preceq\}\mathcal{T}$ and hence the converse.
7. On the failure of the convexity condition $(\omega_{4})$
In this final section we provide the technical construction of a function $\omega:[0,+\infty)\rightarrow[0,+\infty)$ satisfying the following properties:
  • $(I)$
    $\omega$ is normalized, continuous, non-decreasing and $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$.
  • $(II$)
    $\omega$ satisfies $(\omega_{3})$.
  • $(III)$
    $\gamma(\omega)=+\infty$ holds. This implies that $\omega$ also has $(\omega_{\operatorname{snq}})$, $(\omega_{1})$, and that $\omega$ is equivalent to the continuous concave resp. subadditive function $\kappa_{\omega}$; see Lemma 2.6 and Remark 2.5. In fact, we extend (2.10) and even show that $\omega$ is slowly varying, see [2, $(1.2.1)$] and (7.20):
    $\forall\;u>0:\;\;\;\lim_{t\rightarrow+\infty}\frac{\omega(tu)}{\omega(t)}=1.$
  • $(IV)$
    There does not exist any function $\sigma:[0,+\infty)\rightarrow[0,+\infty)$ such that $\sigma$ and $\omega$ are equivalent and such that $\varphi_{\sigma}$ is convex. Thus, in particular, $\varphi_{\omega}$ is not convex. In view of $(I)$ we have that equivalence is described by $O$-growth relations as $t\rightarrow+\infty$; see Section 2.3.
  • $(V)$
    There exist (uncountable) infinitely many pairwise non-equivalent functions satisfying (I)(IV)(I)-(IV).
These properties yield:
  • $(\mathfrak{C}_{1})$
    By (I)(III)(I)-(III), and in view of Remark 2.5 and Lemmas 2.3 and 2.6, we have that $\omega$ is equivalent to a weight function in the sense of Beurling-Björck (see Definition 2.10).
  • $(\mathfrak{C}_{2})$
    By $(IV)$ it follows that in the whole equivalence class of $\omega$ we cannot find any BMT-weight function. In particular, $\omega$ itself is not a BMT-weight function and, indeed, even any arbitrary function being equivalent to $\omega$ violates $(\omega_{4})$.
  • $(\mathfrak{C}_{3})$
    Finally, $(V)$ implies that there exist “quite many” weight functions having this non-standard behavior.
When combining the information from $(\mathfrak{C}_{1})$ and $(\mathfrak{C}_{2})$ with the characterizations obtained in Theorems 4.1 and 4.2 (see also Proposition 4.9), and by taking into account the comments in Section 5.1, especially the characterization of the growth relations in comments $(c)$ and $(d)$ there, within the exponential-type weight function matrix setting we obtain: There exist examples of weighted classes treated in the Beurling-Björck-setting which cannot be described (alternatively) by any BMT-weight function and this statement applies to both Roumieu- and Beurling-type spaces.
This (counter-)example should be compared with Theorem 2.11 and it is somehow converse to this result since we are constructing explicitly a BB-weight and not a BMT-weight function.
7.1. Preparatory results
We collect several preparatory results which are needed in the proof of the construction in the next section.
Lemma 7.1.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be arbitrary. Assume that there exists $\sigma:[0,+\infty)\rightarrow[0,+\infty)$ such that $\omega$ and $\sigma$ are equivalent and such that $\varphi_{\sigma}:t\mapsto\sigma(e^{t})$ is convex (on $\mathbb{R}$), i.e. $\sigma$ has $(\omega_{4})$. Then $\omega$ has to satisfy:
(7.1) $\exists\;A\geq 1\;\forall\;x,y\in\mathbb{R}\;\forall\;0\leq t\leq 1:\;\;\;\varphi_{\omega}(ty+(1-t)x)\leq At\varphi_{\omega}(y)+A(1-t)\varphi_{\omega}(x)+A.$
Proof. The equivalence between $\omega$ and $\sigma$ gives that
$\exists\;B\geq 1\;\forall\;t\geq 0:\;\;\;-1+B^{-1}\sigma(t)\leq\omega(t)\leq B\sigma(t)+B,$
and by taking $s:=e^{t}$ these estimates transfer to the corresponding functions $\varphi_{\omega}$, $\varphi_{\sigma}$ as well (with the same choice for $B$ and for all $t\in\mathbb{R}$).
Thus we can estimate as follows for all $x,y\in\mathbb{R}$ and $0\leq t\leq 1$:
$\varphi_{\omega}(ty+(1-t)x)\leq B\varphi_{\sigma}(ty+(1-t)x)+B\leq Bt\varphi_{\sigma}(y)+B(1-t)\varphi_{\sigma}(x)+B$
$\leq B^{2}t\varphi_{\omega}(y)+B^{2}t+B^{2}(1-t)\varphi_{\omega}(x)+B^{2}(1-t)+B=B^{2}t\varphi_{\omega}(y)+B^{2}(1-t)\varphi_{\omega}(x)+B^{2}+B.$
This verifies (7.1) for the choice $A:=B^{2}+B$. ∎
Lemma 7.2.
Assume that $\omega:[0,+\infty)\rightarrow[0,+\infty)$ satisfies
(7.2) $\liminf_{t\rightarrow+\infty}\omega(t)>1\Leftrightarrow\liminf_{t\rightarrow+\infty}\varphi_{\omega}(t)>1,$
and such that
(7.3) $\forall\;j\in\mathbb{N}_{>0}\;\exists\;0<t_{j}<\frac{1}{j}\;\exists\;a_{j},b_{j}\geq 0:\;\;\;\varphi_{\omega}(t_{j}b_{j}+(1-t_{j})a_{j})=\varphi_{\omega}(b_{j})=\frac{2j\varphi_{\omega}(a_{j})}{1-jt_{j}}.$
Then there does not exist a function $\sigma:[0,+\infty)\rightarrow[0,+\infty)$ such that $\varphi_{\sigma}$ is convex and such that $\omega$ and $\sigma$ are equivalent.
Note that (7.2) is (2.3) and a mild requirement; it is clearly satisfied if $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$.
Proof. First, we claim that
$\exists\;j_{0}\in\mathbb{N}_{>0}\;\forall\;j\geq j_{0}:\;\;\;\varphi_{\omega}(b_{j})>jt_{j}\varphi_{\omega}(b_{j})+j(1-t_{j})\varphi_{\omega}(a_{j})+j.$
For this note that
$\varphi_{\omega}(b_{j})>jt_{j}\varphi_{\omega}(b_{j})+j(1-t_{j})\varphi_{\omega}(a_{j})+j\Leftrightarrow\varphi_{\omega}(b_{j})(1-jt_{j})>j(1-t_{j})\varphi_{\omega}(a_{j})+j$
$\Leftrightarrow 2j\varphi_{\omega}(a_{j})>j(1-t_{j})\varphi_{\omega}(a_{j})+j;$
we have used the second equality in (7.3) and the last estimate is valid since $1-t_{j}<1$ for all $j\in\mathbb{N}_{>0}$ and since $\varphi_{\omega}(a_{j})>1$ for all $j$ sufficiently large which holds by assumption (7.2). By taking into account the first equality in (7.3) we have shown:
$\forall\;C\geq 1\;\exists\;j_{C}\in\mathbb{N}_{>0}\;\forall\;j\geq j_{C}\;\exists\;0<t_{j}<\frac{1}{j}\;\exists\;a_{j},b_{j}\geq 0:$
$\varphi_{\omega}(t_{j}b_{j}+(1-t_{j})a_{j})=\varphi_{\omega}(b_{j})>Ct_{j}\varphi_{\omega}(b_{j})+C(1-t_{j})\varphi_{\omega}(a_{j})+C.$
This violates (7.1) and finishes the proof by Lemma 7.2. ∎
Lemma 7.3.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing and such that $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$. Then $\omega$ satisfies $(P_{\omega,\gamma})$ if and only if
$\exists\;K>1:\;\;\;\limsup_{t\rightarrow+\infty}\frac{\varphi_{\omega}(\gamma\log(K)+t)}{\varphi_{\omega}(t)}<K.$
Proof. We set $s:=e^{t}$ and the definition of $\varphi_{\omega}(t)=\omega(e^{t})$ yields $\frac{\omega(K^{\gamma}s)}{\omega(s)}=\frac{\omega(e^{\gamma\log(K)}e^{t})}{\omega(e^{t})}=\frac{\varphi_{\omega}(\gamma\log(K)+t)}{\varphi_{\omega}(t)}$. ∎
Corollary 7.4.
Let $\omega:[0,+\infty)\rightarrow[0,+\infty)$ be non-decreasing and such that $\lim_{t\rightarrow+\infty}\omega(t)=+\infty$. If
(7.4) $\forall\;\gamma>0:\;\;\;\limsup_{t\rightarrow+\infty}\frac{\varphi_{\omega}(\gamma+t)}{\varphi_{\omega}(t)}\leq 2$
is valid, then $\gamma(\omega)=+\infty$.
Proof. In view of Lemma 7.3 property (7.4) means that $(P_{\omega,\gamma})$ is valid with the (uniform) choice $K:=e$ for any $\gamma>0$. ∎
7.2. Construction of the (counter-)example
We start with the explicit construction of the weight function and involve the information obtained in Section 7.1. The idea is purely geometric and quite technical, therefore we split the proof into several steps.
Step a - Comments on the geometric definition of $\omega$.
We introduce $\omega$ by defining explicitly $\varphi_{\omega}:t\mapsto\omega(e^{t})$ on $0,+\infty)$, i.e. $\omega$ on $1,+\infty)$ via $\omega(s):=\varphi_{\omega}(\log(s))$. The graph of $\varphi_{\omega}$ consists of a set of straight lines and for this we have to introduce three auxiliary sequences $(x_{j})_{j\in\mathbb{N}_{>0}}$, $(\overline{x}_{j})_{j\in\mathbb{N}_{>0}}$ and $(y_{j})_{j\in\mathbb{N}_{>0}}$. We assume that
$\forall\;j\in\mathbb{N}_{>0}:\;\;\;x_{j}<\overline{x}_{j}<y_{j}<x_{j+1},$
and several more growth restrictions; see Step b below. Indeed, $\overline{x}_{j}$ is given as the convex combination of the points $x_{j}$ and $y_{j}$ subject to a value $0<t_{j}<1$ which is coming from the fourth auxiliary sequence $(t_{j})_{j\in\mathbb{N}_{>0}}$.
The graph of $\varphi_{\omega}$ is defined to be
    the straight line connecting the points $(0,0)$ and $(x_{1},\varphi_{\omega}(x_{1}))$ (since we assume $x_{1}>0$),
    the set of straight lines connecting the points $(x_{j},\varphi_{\omega}(x_{j}))$ and $(\overline{x}_{j},\varphi_{\omega}(\overline{x}_{j}))$ resp. $(\overline{x}_{j},\varphi_{\omega}(\overline{x}_{j}))$ and $(y_{j},\varphi_{\omega}(y_{j}))$ resp. $(y_{j},\varphi_{\omega}(y_{j}))$ and $(x_{j+1},\varphi_{\omega}(x_{j+1}))$.
    The slope of the straight line connecting $(x_{j},\varphi_{\omega}(x_{j}))$ and $(\overline{x}_{j},\varphi_{\omega}(\overline{x}_{j}))$ is denoted by $k_{j}$; the slope of the line connecting $(y_{j},\varphi_{\omega}(y_{j}))$ and $(x_{j+1},\varphi_{\omega}(x_{j+1}))$ by $\ell_{j}$.
    Moreover, the slope of the line connecting $(\overline{x}_{j},\varphi_{\omega}(\overline{x}_{j}))$ and $(y_{j},\varphi_{\omega}(y_{j}))$ is identically 0 for all $j\in\mathbb{N}_{>0}$, i.e. $\varphi_{\omega}(\overline{x}_{j})=\varphi_{\omega}(y_{j})$; see (7.11) and (7.13) below.
    This describes $\omega$ on $1,+\infty)$ with $\omega(1)=\varphi_{\omega}(0)=0$ and in order to complete the definition we set $\omega(t):=0$ for $0\leq t<1$ (i.e. normalization for $\omega$).
Finally, for treating $(V)$ we introduce the following notion: A sequence $\delta=(\delta_{j})_{j\in\mathbb{N}_{>0}}\in\mathbb{R}_{>0}^{\mathbb{N}}$ is called an admissible parameter sequence if
(7.5) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\delta_{j+1}\leq\delta_{j}\leq(j+2)\delta_{j+1},\hskip 25.0pt\lim_{j\rightarrow+\infty}\delta_{j}j=+\infty.$
An immediate example for such a sequence is $\delta_{j}:=\frac{1}{\log(e+j)}$.
Let $\delta=(\delta_{j})_{j\in\mathbb{N}_{>0}}$ be an admissible parameter sequence, arbitrary but from now on fixed. The definition of $\omega$ is depending on chosen $\delta$ which suggests the notation $\varphi_{\omega_{\delta}}$. However, since $\delta$ is now fixed, in order to lighten notation we skip $\delta$ and simply write $\varphi_{\omega}$ except in the very last part Step g: There we verify $(V)$ and hence we have to emphasize the different parameter sequences.
Step b - Definition of the auxiliary sequence $(t_{j})_{j\in\mathbb{N}_{>0}}$.
We choose $0<t_{1}<1$ arbitrary and then set
(7.6) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;t_{j+1}:=\frac{1}{2}\left(\frac{1}{j}-t_{j}\right).$
Let us show by induction that this sequence satisfies
(7.7) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;0<t_{j}<\frac{1}{j},$
which should be compared with (7.3). The case $j=1$ is obvious and note that $t_{j+1}>0$ if and only if $t_{j}<\frac{1}{j}$ is verified, see (7.6). Moreover,
$t_{j+1}<\frac{1}{j+1}\Leftrightarrow\frac{1}{j}-t_{j}<\frac{2}{j+1}\Leftrightarrow\frac{1-j}{j(j+1)}=\frac{1}{j}-\frac{2}{j+1}<t_{j},$
which clearly holds since $t_{j}>0$ by induction hypothesis. From now on let such a sequence $(t_{j})_{j\in\mathbb{N}_{>0}}$ be given and fixed and note that this sequence is not depending on the chosen admissible parameter sequence $\delta$.
Step c - Definition of the required auxiliary sequences and basic growth properties.
We introduce the required auxiliary sequences $(x_{j})_{j\in\mathbb{N}_{>0}}$, $(\overline{x}_{j})_{j\in\mathbb{N}_{>0}}$ and $(y_{j})_{j\in\mathbb{N}_{>0}}$, collect the required growth properties and define $\varphi_{\omega}$ on these values.
Let us set
(7.8) $x_{1}:=\frac{2}{t_{1}},$
and when $x_{j}$, $j\in\mathbb{N}_{>0}$, is given, then define
(7.9) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\varphi_{\omega}(x_{j}):=\delta_{j}jx_{j},$
(7.10) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;y_{j}:=\frac{2\delta_{j}^{-1}\varphi_{\omega}(x_{j})}{1-jt_{j}}=\frac{2jx_{j}}{1-jt_{j}},$
(7.11) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\varphi_{\omega}(y_{j}):=\frac{2j\varphi_{\omega}(x_{j})}{1-jt_{j}}=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}},$
(7.12) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\overline{x}_{j}:=t_{j}y_{j}+(1-t_{j})x_{j},$
(7.13) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\varphi_{\omega}(\overline{x}_{j}):=\varphi_{\omega}(y_{j})\left(=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}}\right),$
(7.14) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;x_{j+1}:=(j+1)y_{j}=\frac{2j(j+1)x_{j}}{1-jt_{j}}.$
Concerning the denominators in the defining expressions recall that by (7.7) we get $0<1-jt_{j}<1$ for any $j\in\mathbb{N}_{>0}$. Moreover, we immediately have $1<\frac{2}{t_{1}}=x_{1}$, that $x_{j}<\overline{x}_{j}<y_{j}<x_{j+1}$ for all $j\in\mathbb{N}_{>0}$ and that all three sequences tend to infinity as $j\rightarrow+\infty$. By (7.9) the slope of the straight line connecting $(0,0)$ and $(x_{1},\varphi_{\omega}(x_{1}))$ is equal to $\delta_{1}$. Next we prove
(7.15) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;x_{j}\geq\frac{2j}{t_{j}}(>2j^{2}\geq 2j).$
The case $j=1$ is precisely (7.8). By (7.14) we get
$x_{j+1}\geq\frac{2(j+1)}{t_{j+1}}\Leftrightarrow\frac{2j(j+1)x_{j}}{1-jt_{j}}\geq\frac{2(j+1)}{t_{j+1}}\Leftrightarrow\frac{jx_{j}}{1-jt_{j}}\geq\frac{1}{t_{j+1}},$
and (7.6) yields
$\frac{1}{t_{j+1}}\leq\frac{2j^{2}}{t_{j}(1-jt_{j})}\Leftrightarrow\frac{t_{j}(1-jt_{j})}{2j^{2}}\leq t_{j+1}\Leftrightarrow\frac{t_{j}}{2j}\left(\frac{1}{j}-t_{j}\right)\leq\frac{1}{2}\left(\frac{1}{j}-t_{j}\right),$
which is clearly satisfied (recall that $t_{j}<1$). Finally, the induction hypothesis implies $\frac{jx_{j}}{1-jt_{j}}\geq\frac{2j^{2}}{t_{j}(1-jt_{j})}$ and by combining everything we have shown the desired estimate for $j+1$, thus (7.15) is verified. Next we study the growth differences between the auxiliary sequences; these identities are then crucial for Step f below.
Claim: One has $\min\{y_{j}-\overline{x}_{j},y_{j}-x_{j},x_{j+1}-y_{j},\overline{x}_{j}-x_{j}\}\geq j$ for all $j\in\mathbb{N}_{>0}$.
First note that, since $x_{j}<\overline{x}_{j}$, for having $y_{j}-x_{j}\geq j$ it suffices to verify $y_{j}-\overline{x}_{j}\geq j$.
Concerning this estimate, by taking into account (7.12) and (7.10) we get
$y_{j}-\overline{x}_{j}=(1-t_{j})y_{j}-(1-t_{j})x_{j}=(1-t_{j})(y_{j}-x_{j})=(1-t_{j})\left(\frac{2jx_{j}}{1-jt_{j}}-x_{j}\right)\geq j$
$\Leftrightarrow jx_{j}\geq j\frac{1-jt_{j}}{2(1-t_{j})}+x_{j}\frac{1-jt_{j}}{2}.$
This last estimate holds since $\frac{1-jt_{j}}{2}\leq\frac{1}{2}$ and $\frac{1-jt_{j}}{2(1-t_{j})}\leq\frac{1}{2}$, so we get $j\frac{1-jt_{j}}{2(1-t_{j})}+x_{j}\frac{1-jt_{j}}{2}\leq\frac{j+x_{j}}{2}$ and the estimate $j+x_{j}\leq 2jx_{j}$ is clear because $x_{j}\geq j$ for any $j\in\mathbb{N}_{>0}$ (see (7.15)).
Next, we see that (7.14) yields $x_{j+1}-y_{j}=jy_{j}>jx_{j}\geq j$.
Finally, by (7.12) and (7.10) we get
$\overline{x}_{j}-x_{j}=t_{j}(y_{j}-x_{j})=t_{j}\left(\frac{2jx_{j}}{1-jt_{j}}-x_{j}\right)\geq j\Leftrightarrow jx_{j}\geq j\frac{1-jt_{j}}{2t_{j}}+x_{j}\frac{1-jt_{j}}{2},$
and let us see that this estimate holds: First, $x_{j}\frac{1-jt_{j}}{2}<\frac{x_{j}}{2}$ is clear and so it suffices to have $\left(j-\frac{1}{2}\right)x_{j}\geq j\frac{1-jt_{j}}{2t_{j}}$. Second, by taking into account (7.15) for this estimate it suffices to have $\left(j-\frac{1}{2}\right)\frac{2j}{t_{j}}\geq\frac{j}{2t_{j}}$ which is equivalent to having $4j^{2}-2j\geq j$, i.e. to $\frac{4}{3}j\geq 1$ which is clear for all $j\in\mathbb{N}_{>0}$. Thus the claim is verified.
Now let us compute and compare the slopes $k_{j}$ and $\ell_{j}$ introduced in Step a explicitly. By definition we have
$k_{j}$ $:=\frac{\varphi_{\omega}(\overline{x}_{j})-\varphi_{\omega}(x_{j})}{\overline{x}_{j}-x_{j}}=\delta_{j}\frac{2j^{2}x_{j}(1-jt_{j})^{-1}-jx_{j}}{t_{j}(y_{j}-x_{j})}=\delta_{j}\frac{2j^{2}x_{j}-jx_{j}(1-jt_{j})}{t_{j}(2jx_{j}(1-jt_{j})^{-1}-x_{j})(1-jt_{j})}$
$=\delta_{j}\frac{jx_{j}(2j-1+jt_{j})}{t_{j}(2jx_{j}-x_{j}+jx_{j}t_{j})}=\delta_{j}\frac{jx_{j}(2j-1+jt_{j})}{t_{j}x_{j}(2j-1+jt_{j})},$
and so
(7.16) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;k_{j}=\frac{\delta_{j}j}{t_{j}}.$
Since $k_{j}=\frac{\delta_{j}j}{t_{j}}>\delta_{j}j^{2}$ (see (7.7)), the second part of (7.5) implies $\lim_{j\rightarrow+\infty}k_{j}=+\infty$.
Similarly, for $\ell_{j}$ one has by definition that
$\ell_{j}$ $:=\frac{\varphi_{\omega}(x_{j+1})-\varphi_{\omega}(y_{j})}{x_{j+1}-y_{j}}=\frac{2\delta_{j+1}j(j+1)^{2}x_{j}(1-jt_{j})^{-1}-2\delta_{j}j^{2}x_{j}(1-jt_{j})^{-1}}{2j(j+1)x_{j}(1-jt_{j})^{-1}-2jx_{j}(1-jt_{j})^{-1}}$
$=\frac{\delta_{j+1}(j+1)^{2}-\delta_{j}j}{(j+1)-1}=\frac{\delta_{j+1}(j+1)^{2}-\delta_{j}j}{j}.$
The first part in (7.5) implies, on the one hand, that $\ell_{j}\leq\frac{\delta_{j}(j^{2}+j+1)}{j}$ because $\delta_{j+1}\leq\delta_{j}$, and, on the other hand, also $\ell_{j}>0\Leftrightarrow\delta_{j+1}\frac{(j+1)^{2}}{j}>\delta_{j}$ holds since $\frac{(j+1)^{2}}{j}=j+2+\frac{1}{j}>j+2$. Consequently, we have verified
(7.17) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;0<\ell_{j}\leq\frac{\delta_{j}(j^{2}+j+1)}{j}.$
Moreover, it holds that
(7.18) $\forall\;j\in\mathbb{N}_{>0}:\;\;\;\ell_{j}<k_{j+1}$
because, first, by (7.16) we get
$k_{j+1}>\ell_{j}\Leftrightarrow\frac{\delta_{j+1}(j+1)}{t_{j+1}}>\frac{\delta_{j+1}(j+1)^{2}-\delta_{j}j}{j}\Leftrightarrow\delta_{j+1}(j+1)>t_{j+1}\left(\delta_{j+1}\left(j+2+\frac{1}{j}\right)-\delta_{j}\right).$
This estimate is valid since
$t_{j+1}\left(\delta_{j+1}\left(j+2+\frac{1}{j}\right)-\delta_{j}\right)<t_{j+1}\delta_{j+1}\left(j+2+\frac{1}{j}\right)<\frac{1}{j+1}\delta_{j+1}\left(j+2+\frac{1}{j}\right),$
which follows by taking into account (7.7) and then, in order to conclude, note that $\frac{1}{j+1}\left(j+2+\frac{1}{j}\right)<(j+1)$ is obvious.
Summarizing, from the given construction it follows that $\varphi_{\omega}:[0,+\infty)\rightarrow[0,+\infty)$ is continuous and non-decreasing with $\lim_{t\rightarrow+\infty}\varphi_{\omega}(t)=+\infty$. The last property follows, in particular, also by the next step.
Step d - $\omega$ satisfies $(\omega_{3})$ (and hence $(\omega_{3,w})$, too).
In order to conclude we have to show that $\lim_{t\rightarrow+\infty}\frac{\varphi_{\omega}(t)}{t}=+\infty$ and by construction and definition it suffices to investigate the “corner points” of the graph; i.e. that
(7.19) $\lim_{j\rightarrow+\infty}\frac{\varphi_{\omega}(x_{j})}{x_{j}}=\lim_{j\rightarrow+\infty}\frac{\varphi_{\omega}(y_{j})}{y_{j}}=+\infty.$
For this note that $x\mapsto\frac{kx+d}{x}=k+\frac{d}{x}$ is non-increasing on $(0,+\infty)$ if and only if $d\geq 0$ and in view of the geometric construction recall that the slopes of the lines connecting $(\overline{x}_{j},\varphi_{\omega}(\overline{x}_{j}))$ and $(y_{j},\varphi_{\omega}(y_{j}))$ are identically 0.
For all $j\in\mathbb{N}_{>0}$ by (7.9) we get
$\frac{\varphi_{\omega}(x_{j})}{x_{j}}=\frac{\delta_{j}jx_{j}}{x_{j}}=\delta_{j}j,$
and (7.10) and (7.11) together give
$\frac{\varphi_{\omega}(y_{j})}{y_{j}}=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}}\frac{1-jt_{j}}{2jx_{j}}=\delta_{j}j,$
which proves (7.19) by taking into account the second part in (7.5).
[…]
We apply Lemma 7.2. First note that as a consequence of Step c and Step d we have that $\lim_{t\rightarrow+\infty}\varphi_{\omega}(t)=+\infty$ and hence, in particular, (7.2) is valid. In order to conclude we prove (7.3) and verify this condition for the choices $t_{j}$ from Step b (recall (7.7)) and when taking $a_{j}:=x_{j}$, $b_{j}:=y_{j}$. By combining (7.12), (7.13) and (7.9) we get for all $j\in\mathbb{N}_{>0}$
$\varphi_{\omega}(t_{j}y_{j}+(1-t_{j})x_{j})=\varphi_{\omega}(\overline{x}_{j})=\varphi_{\omega}(y_{j})=\frac{2\delta_{j}j^{2}x_{j}}{1-jt_{j}}=\frac{2j\varphi_{\omega}(x_{j})}{1-jt_{j}},$
which is precisely (7.3) for the aforementioned choices.
Step f - $\omega$ satisfies $\gamma(\omega)=+\infty$.
In order to prove $\gamma(\omega)=+\infty$ we want to apply Corollary 7.4 and so the aim is to verify (7.4). In fact we even show
(7.20) $\forall\;\gamma>0:\;\;\;\lim_{s\rightarrow+\infty}\frac{\varphi_{\omega}(\gamma+s)}{\varphi_{\omega}(s)}=1\Leftrightarrow\lim_{t\rightarrow+\infty}\frac{\omega(e^{\gamma}t)}{\omega(t)}=1,$
i.e. $\omega$ is slowly varying, see [2, $(1.2.1)$], [12, Sect. 2.2] and the comments at the end of Section 2.3.
Let $\gamma>0$ be given, arbitrary large but from now on fixed. By the previous Claim in Step c it follows that for all sufficiently large $s\in[0,+\infty)$ (depending on given $\gamma$) we eventually have to distinguish between the following three different cases.
Case i - Let $x_{j}\leq s\leq\overline{x}_{j}$, for $j\in\mathbb{N}_{>0}$ sufficiently large depending on $\gamma$, then by the claim $\gamma+s\leq y_{j}$.
In this situation, by definition of the graph and the slope $k_{j}$ and since $\varphi_{\omega}$ is non-decreasing, we have the following estimate:
$\frac{\varphi_{\omega}(\gamma+s)}{\varphi_{\omega}(s)}\leq\frac{\varphi_{\omega}(s)}{\varphi_{\omega}(s)}+\frac{\gamma k_{j}}{\varphi_{\omega}(s)}\leq 1+\frac{\gamma k_{j}}{\varphi_{\omega}(x_{j})}.$
(7.9), (7.16) and (7.15) together imply
$\frac{k_{j}}{\varphi_{\omega}(x_{j})}=\frac{\delta_{j}j}{t_{j}\delta_{j}jx_{j}}=\frac{1}{t_{j}x_{j}}\leq\frac{1}{2j},$
which tends to 0 as $j\rightarrow+\infty$.
Case ii - Let $\overline{x}_{j}\leq s\leq y_{j}$, then the claim gives that $\gamma+s\leq x_{j+1}$ for all $j\in\mathbb{N}_{>0}$ sufficiently large.
In this situation, by definition of the graph and the slope $\ell_{j}$ and since $\varphi_{\omega}$ is non-decreasing, we have the following estimate:
$\frac{\varphi_{\omega}(\gamma+s)}{\varphi_{\omega}(s)}\leq\frac{\varphi_{\omega}(s)}{\varphi_{\omega}(s)}+\frac{\gamma\ell_{j}}{\varphi_{\omega}(s)}\leq 1+\frac{\gamma\ell_{j}}{\varphi_{\omega}(\overline{x}_{j})}.$
(7.13), (7.17) and (7.15) together with (7.7) imply
$\frac{\ell_{j}}{\varphi_{\omega}(\overline{x}_{j})}\leq\frac{\delta_{j}(j^{2}+j+1)}{j2j^{2}\delta_{j}x_{j}(1-jt_{j})^{-1}}\leq\frac{(j^{2}+j+1)(1-jt_{j})t_{j}}{4j^{4}}\leq\frac{j^{2}+j+1}{4j^{4}},$
which tends to 0 as $j\rightarrow+\infty$.
Case iii - Let $y_{j}\leq s\leq x_{j+1}$, then $\gamma+s\leq\overline{x}_{j+1}$ for all $j\in\mathbb{N}_{>0}$ sufficiently large.
In this situation we have, by taking into account (7.18), that
$\frac{\varphi_{\omega}(\gamma+s)}{\varphi_{\omega}(s)}\leq\frac{\varphi_{\omega}(s)}{\varphi_{\omega}(s)}+\frac{\gamma\max\{\ell_{j},k_{j+1}\}}{\varphi_{\omega}(s)}\leq 1+\frac{\gamma k_{j+1}}{\varphi_{\omega}(y_{j})}.$
Then note that by (7.6) and the fact that $t_{j}<1$ for all $j\in\mathbb{N}_{>0}$ we have $\frac{(1-jt_{j})t_{j}}{t_{j+1}}\leq 2j\Leftrightarrow(1-jt_{j})t_{j}\leq(1-jt_{j})$. By combining this with (7.11), (7.16), (7.15) and property $\delta_{j+1}\leq\delta_{j}$ (see the first part in (7.5)) we obtain
$\frac{k_{j+1}}{\varphi_{\omega}(y_{j})}=\frac{\delta_{j+1}(j+1)}{t_{j+1}2\delta_{j}j^{2}x_{j}(1-jt_{j})^{-1}}\leq\frac{(j+1)(1-jt_{j})t_{j}}{t_{j+1}4j^{3}}\leq\frac{j+1}{2j^{2}},$
which tends to 0 as $j\rightarrow+\infty$.
Summarizing, Cases i-iii verify (7.20) and hence this part is finished.
Step g - $\omega$ satisfies $(\omega_{\operatorname{nq}})$ and $(\alpha_{0})$; more precisely, $\omega$ is equivalent to $\kappa_{\omega}$ (see (2.9)).
The fact that $\omega$ has $(\omega_{\operatorname{nq}})$ follows from Step f and (2.10) whereas $(\alpha_{0})$ holds by Step f and (2.11). By Remark 2.5 the function $\omega$ is equivalent to $\kappa_{\omega}$ which is continuous, sub-additive and $\kappa_{\omega}(0)=\omega(0)=0$.
Step h - There exist (uncountable) infinitely many pairwise non-equivalent weights satisfying all the requirements shown above.
Indeed, we are going to verify that if $\delta=(\delta_{j})_{j\in\mathbb{N}_{>0}}$ and $\delta^{\prime}=(\delta^{\prime}_{j})_{j\in\mathbb{N}_{>0}}$ are two admissible parameter sequences satisfying
(7.21) $\sup_{j\in\mathbb{N}_{>0}}\frac{\delta_{j}}{\delta^{\prime}_{j}}=+\infty\;\;\;\text{or}\;\;\;\sup_{j\in\mathbb{N}_{>0}}\frac{\delta^{\prime}_{j}}{\delta_{j}}=+\infty,$
then the corresponding weights $\omega_{\delta}$ and $\omega_{\delta^{\prime}}$ are not equivalent.
For this apply (7.9) and when assuming the first condition in (7.21) we get
$\sup_{x\geq 0}\frac{\varphi_{\omega_{\delta}}(x)}{\varphi_{\omega_{\delta^{\prime}}}(x)}\geq\sup_{j\in\mathbb{N}_{>0}}\frac{\varphi_{\omega_{\delta}}(x_{j})}{\varphi_{\omega_{\delta^{\prime}}}(x_{j})}=\sup_{j\in\mathbb{N}_{>0}}\frac{\delta_{j}jx_{j}}{\delta^{\prime}_{j}jx_{j}}=\sup_{j\in\mathbb{N}_{>0}}\frac{\delta_{j}}{\delta^{\prime}_{j}}=+\infty.$
[…]
This finishes the construction of the desired (counter-)example resp. of the uncountable infinite family of (counter-)examples.
Remark 7.5.
Let an admissible parameter sequence $\delta$ satisfying $\lim_{j\rightarrow+\infty}\delta_{j}=0$ be given; consider e.g. $\delta_{j}:=\frac{1}{\log(e+j)}$. However, note that this additional condition is not required necessarily in the above arguments and $\delta$ can be bounded by below by some $\epsilon>0$. In any case, under this assumption each power sequence $\delta^{\alpha}=(\delta_{j}^{\alpha})_{j\in\mathbb{N}_{>0}}$, $0<\alpha<1$, is again an admissible parameter sequence since $\delta_{j+1}^{\alpha}\leq\delta_{j}^{\alpha}\leq(j+2)^{\alpha}\delta_{j+1}^{\alpha}\leq(j+2)\delta_{j+1}^{\alpha}$ for all $j\in\mathbb{N}_{>0}$ and $\delta_{j}^{\alpha}\geq\delta_{j}$ for all $j$ such that $\delta_{j}\leq 1$; i.e. for all $j$ sufficiently large. Thus all requirements from (7.5) hold for $\delta^{\alpha}:=(\delta^{\alpha}_{j})_{j\in\mathbb{N}_{>0}}$ and (each) $\delta^{\prime}:=\delta^{\alpha}$ serves as an example for an admissible parameter sequence satisfying the second condition in (7.21).
Note that in general it is not clear that $\delta^{\alpha}$ is admissible if $\alpha>1$.
Remark 7.6.
Let $\mathbf{M}=(M_{p})_{p\in\mathbb{N}}\in\mathbb{R}_{>0}^{\mathbb{N}}$ be given, then the associated function $\omega_{\mathbf{M}}:[0,+\infty)\rightarrow[0,+\infty)\cup\{+\infty\}$ is defined as follows:
(7.22) $\omega_{\mathbf{M}}(t):=\sup_{p\in\mathbb{N}}\log\left(\frac{M_{0}t^{p}}{M_{p}}\right),\qquad t\geq 0,$
with the conventions $0^{0}:=1$ and $\log(0)=$-$\infty$. When treating this function, naturally one should assume that $\lim_{p\rightarrow+\infty}(M_{p})^{1/p}=+\infty$ holds in order to ensure $\omega_{\mathbf{M}}:[0,+\infty)\rightarrow[0,+\infty)$ (well-definedness). For more details we refer to [16, Chapitre I], the more recent work [33] and see also [15, Sect. 3]. It is known and immediate by definition that $(\omega_{4})$ holds for $\omega_{\mathbf{M}}$ and consequently in the whole equivalence class (w.r.t. relation $\sim$) of the previously constructed function there does not exist any associated weight function. In particular, the obtained (counter-)example cannot be described within the weight sequence setting which means that $\omega\equiv\omega_{\mathbf{M}}$ fails for any weight sequence $\mathbf{M}$.
On the other hand the above constructed $\omega$ is matrix admissible since crucially $(\omega_{3})$ holds.
We finish by mentioning the following question/goal: Transfer the construction of $\varphi_{\omega}$ (resp. of $\omega$) in Section 7 to the higher dimensional case (anisotropic setting). We expect that the analogous statement holds as well but such an explicit construction is more technical and involved and also the arguments using the growth index $\gamma(\cdot)$ are becoming unclear in this general situation since so far this quantity is only defined and studied for $\omega:[0,+\infty)\rightarrow[0,+\infty)$.